### v1

# Extended Resolution Clause Learning via Dual Implication Points

Abstract
We present a new extended resolution clause learning (ERCL) algorithm, implemented as part of a conflict-driven clause-learning (CDCL) SAT solver, wherein new variables are dynamically introduced as definitions for Dual Implication Points (DIPs) in the implication graph constructed by the solver at runtime. DIPs are generalizations of unique implication points and can be informally viewed as a pair of dominator nodes, from the decision variable at the highest decision level to the conflict node, in an implication graph. We perform extensive experimental evaluation to establish the efficacy of our ERCL method, implemented as part of the MapleLCM SAT solver and dubbed xMapleLCM, against several leading solvers including the baseline MapleLCM, as well as CDCL solvers such as Kissat 3.1.1, CryptoMiniSat 5.11, and SBVA+CaDiCaL, the winner of SAT Competition 2023. We show that xMapleLCM outperforms these solvers on Tseitin and XORified formulas. We further compare xMapleLCM with GlucoseER, a system that implements extended resolution in a different way, and provide a detailed comparative analysis of their performance.
keywords:
SAT, Extended resolution, Conflict Analysis
category:
[…]
1 Introduction
Over the last several years, Conflict-Driven Clause-Learning (CDCL) SAT solvers have had a dramatic impact on many fields including software engineering [23], security [39], and AI [22, 21]. As solvers continue to be adopted in increasing complex settings, the demand for greater efficiency and reasoning power by users continues unabated.
While developers continue to improve CDCL SAT solvers, it is simultaneously true that these solvers are provably no more powerful than the relatively weak general resolution (Res) proof system [1, 32], and therefore are fundamentally limited. Hence, solver developers have been actively researching novel algorithms that implement stronger proof systems that go beyond Res. Examples of such algorithms include satisfaction-driven clause-learning (SDCL) SAT solvers [17, 31], bounded variable addition (BVA) [26], symmetry breaking [35], and extended resolution (ER) solvers such as GlucoseER [3].
Continuing this trend of strong proof system implementations, we present a new extended resolution clause learning (ERCL) algorithm, incorporated into a CDCL SAT solver, where new variables are dynamically introduced as definitions for dual implication points (DIPs) in the implication graph constructed by the solver at run time. The concept of a DIP is best understood as a generalization of unique implication points (UIPs). Informally, a UIP can be defined as a dominator node in an implication graph, corresponding to a variable at the highest decision level (DL), that dominates all paths from the decision variable node at the highest DL to the conflict node. By contrast, a DIP is a pair of dominator nodes in an implication graph such that any path from the decision variable node at the highest DL to the conflict node must pass through at least one node in the pair [^1] While it is natural to generalize the concept of a DIP to K-Implication Points or $k$-IPs, we do not discuss them in this paper..
Implementation of the ERCL algorithm requires several additional methods. First, we need a method to identify DIPs, i.e., a technique that takes as input an implication graph and outputs a DIP and does so in time linear in the size of the input. Second, we need a technique replaces this DIP pair with a new variable and appropriately modifies the clause learning algorithm to learn new clauses involving DIPs. Third, we need an ER framework, built on top of a CDCL SAT solver, that enables new variable addition, ER clause addition and deletion, etc. Finally, we need heuristics that specialize the above mentioned methods in a variety of ways, such as clause learning and clause deletion policies that are based on different kinds of DIPs. We implement all of these methods as part of the MapleLCM solver [25], and refer to the resulting solver as xMapleLCM. In fact our proposed ERCL method, and its implementation, is very general and easily extensible thus encouraging future exploration and specialization efforts with a variety of heuristics.
Contributions.
  1. 1.
    DIP: We introduce the concept of dual implication points (DIP), a generalization of UIPs in conflict graphs. We also came up with an algorithm that computes them in linear time. However, due to space limitations and to the non-trivial nature of the procedure, we discuss this in a separate paper [11].
  2. 2.
    ERCL Algorithm: We introduce a highly parameterizable DIP-based ERCL algorithm. The existence of a multitude of different DIPs in a single conflict graph allows us to derive a large variety of ERCL algorithms. This flexibility is crucial in adapting the procedure to different scenarios, unlike previous methods that couple CDCL with extended resolution.
  3. 3.
    xMapleLCM: We present a highly extensible and general ER framework as part of xMapleLCM, which allows developers to easily add their own new variable addition, ER clause learning/deletion, and branching policies. Given that DIP-based ERCL is highly flexible by nature, such a framework is necessary to quickly prototype new procedures.
  4. 4.
    Experiments: We perform extensive empirical evaluation and ablation studies on four different classes of instances, namely, SAT Competition 2023 Main Track, random-k-XOR, Tseitin, and interval matching, and compare xMapleLCM against leading solvers such as Kissat 3.1.1 [8], CryptoMiniSat 5.11 [36], SBVA CaDiCaL [14], and the extended-resolution solver GlucoseER [4]. Results show that on the last three sets of hard combinatorial formulas, CDCL SAT solvers perform very poorly, whereas both xMapleLCM and GlucoseER excel on them. Hence, we can now finally state that extended resolution can be added to CDCL and improve the performance of these solvers, and that there are at least two completely different ways to achieve this. Moreover, a simple heuristic dynamically allows us to detect whether extended resolution is being helpful and turn back to standard CDCL in order to get the best of both worlds. This technique enables xMapleLCM to perform similarly to MapleLCM on the SAT Competition 2023 Main Track instances.
2 Related Work
The idea of using ER in SAT solving has been studied in various forms in the literature for nearly two decades. The closest approach to ours is GlucoseER [3], where extended variables are introduced dynamically during the CDCL search: whenever two consecutive learned lemmas are of the form $\lnot l_{1}\lor C$ and $\lnot l_{2}\lor C$, with $l_{1}$ and $l_{2}$ being their UIPs, then an extended variable $z\leftrightarrow l_{1}\lor l_{2}$ is generated and any future lemma of the form $l_{1}\lor l_{2}\lor D$ is replaced by $z\lor D$.
Our method differs significantly from GlucoseER in the way extended variables are identified: we choose DIPs, which can be seen as pairs of variables for which adding a definition would create a better first UIP (often written as 1UIP) in the conflict graph, whereas GlucoseER definitions are constructed using already existing UIPs. Also, unlike in GlucoseER, our approach does not always learn the standard 1UIP clause, but multiple clauses might be learned that take into account the newly introduced variable. On the other hand, there are similarities at the heart of both procedures: a certain restriction of ER is considered and introduced variables are used to shorten subsequently found clauses.
More recent uses of ER in SAT solvers are Bounded Variable Addition (BVA) [26] and its structured version SBVA [14], that due to clever strategies are able to identify new extended variables whose introduction can reduce the number of clauses. Essentially, the ultimate goal of BVA techniques is to reduce the size of the formula by reducing the number of clauses at the cost of adding a new variable. One big difference with our work is that this process is done only in a preprocessing step. Finally, one additional direction that has been researched is the development of a BDD-based solver to generate ER proofs [10].
Other approaches aiming at improving CDCL solvers by allowing them to use a more powerful proof system are related to the Propagation Redundancy notion [16, 18], either via preprocessing steps [34] or via the use of the SDCL algorithm [19, 31]. While these methods implement proof systems that are stronger than Res, many of them (without new variable addition) are known to be weaker than ER.
3 Preliminaries
We assume that the reader is familiar with the satisfiability (SAT) problem and the CDCL algorithm, and we refer her to the Handbook of Satisfiability for an excellent overview of these topics [27]. Below we focus on conflict analysis, which is the most relevant ingredient from the CDCL algorithm for this paper. We do so by means of the following example.
Example 3.1.
Consider the following clauses
$\begin{array}[]{llllll}(1)&y_{1}\lor\lnot x_{1}\lor x_{2}&(6)&\lnot x_{5}\lor x_{7}&(11)&\lnot x_{11}\lor x_{12}\\ (2)&\lnot x_{1}\lor\lnot x_{3}&(7)&x_{6}\lor\lnot x_{7}\lor x_{8}&(12)&x_{10}\lor\lnot x_{11}\lor x_{13}\\ (3)&y_{2}\lor\lnot x_{1}\lor x_{4}&(8)&\lnot y_{3}\lor\lnot y_{4}\lor\lnot x_{5}\lor\lnot x_{9}&(13)&x_{12}\lor\lnot x_{13}\\ (4)&\lnot y_{3}\lor\lnot x_{2}\lor x_{3}\lor\lnot x_{4}\lor x_{5}&(9)&\lnot y_{4}\lor x_{9}\lor\lnot x_{10}\\ (5)&y_{1}\lor\lnot x_{5}\lor\lnot x_{6}&(10)&\lnot y_{5}\lor y_{6}\lor\lnot x_{8}\lor x_{9}\lor x_{11}\\ \end{array}$
Assume that CDCL has constructed an assignment that contains, among others, literals $\{\lnot y_{1},\lnot y_{2},y_{3},y_{4},y_{5},\lnot y_{6}\}$. Since no propagation is possible, it now decides to add the decision literal $x_{1}$. Due to clause (1), we can unit propagate literal $x_{2}$, being (1) the reason of $x_{2}$ and its antecedents $\{\lnot y_{1},x_{1}\}$. Similarly, $\lnot x_{3}$ is propagated due to reason (2), with antecedents $\{x_{1}\}$, and $x_{4}$ due to reason (3) with antecedents $\{\lnot y_{2},x_{1}\}$. If we continue this process we eventually find that clause (13) is conflicting and we can construct the conflict graph in Figure 1, where every literal in the current decision level has incoming edges corresponding to its antecedents (except of course the decision literal). A special conflict node $\perp$ with incoming edges $\{\overline{x_{12}},x_{13}\}$ represents conflicting clause (13).
The graph clearly shows that if we set $x_{1}$ to true, together with the literals of previous decision levels (the $y$’s), we obtain a conflict. However, the same happens with $x_{5}$, since any path from $x_{1}$ to the conflict necessarily goes through $x_{5}$. Literals with this property are called Unique Implication Points (UIPs), of which we only have $x_{1}$ and $x_{5}$. Since $x_{5}$ is the one closest to the conflict we call it First Unique Implication Point (1UIP) [40]. It is easy to see that if we set $x_{5}$ and the $y$ literals that enter the cut delimited by the blue line, unit propagation derives the same conflict. Hence, since they cannot be simultaneously true, we can learn $\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor x_{5}$. The quality of a lemma can be assessed by its Literal Block Distance (LBD) [4]: the number of different decision levels of the literals in the lemma. The lower the LBD, the better the lemma. In our case, if we are at decision level 5, $y_{4}$ and $y_{6}$ belong to decision level 2, and $y_{5}$ to decision level 4, the LBD of the lemma is 3. $\Box$
$x_{1}$$x_{2}$$\overline{x_{3}}$$x_{4}$$\overline{y_{2}}$$\overline{y_{1}}$$x_{5}$$y_{3}$$\overline{x_{6}}$$x_{7}$$\overline{x_{9}}$$y_{5}$$x_{8}$$\overline{y_{6}}$$x_{11}$$\overline{x_{10}}$$y_{4}$$\overline{x_{12}}$$x_{13}$$\perp$
Figure 1: Conflict graph associated with Example 3.1. If 1UIP learning is applied, we generate lemma $\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor x_{5}$. White nodes belong to the current decision level, whereas blue ones are from previous decision levels.
Resolution. Given two clauses $l\lor C$ and $\lnot l\lor D$, the resolution inference rule allows one to derive the logical consequence $C\lor D$. It is well known that the lemma derived in Example 3.1 can be obtained via a series of resolution steps that start with the conflicting clause, and resolve with reasons of literals in the reverse order in which they were added to the assignment. In fact, it has been proved [33, 2] that resolution and CDCL (with restarts) are polynomially equivalent, and hence classes of formulas that are hard for resolution; e.g., the pigeonhole principle - PHP [15], or Tseitin formulas [38]) are also hard for CDCL SAT solvers.
Extended Resolution (ER). Given two literals $l_{1}$ and $l_{2}$, the extended resolution [37] rule allows us to introduce clauses representing the definition $z\leftrightarrow l_{1}\lor l_{2}$. ER can be substantially more powerful than resolution; for instance, it allows polynomial size proofs of PHP [13]. Incorporating ER to CDCL solvers could potentially enable them to solve such formulas in polynomial time. However, we lack good methods to incorporate ER into CDCL proof search. This is precisely the aim of this paper, namely, incorporating a restricted version of ER into CDCL.
4 Dual Implication Points
As discussed in the previous section, UIPs are crucial for conflict clause learning in the CDCL algorithm. We now introduce a new concept of a Dual Implication Point (DIP) that gives a tool for analyzing the conflict graph. Its applications include discovering new implied 2-clauses, introducing new variables by extension, and learning clauses involving the extension variables. The idea behind a dual implication point is that it consists of a pair of vertices (literals) in the conflict graph that disconnects or “separates” the decision literal from the contradiction. [^2] Perhaps “Dual Implication Pair” would be a better name than “Dual Implication Point’, since a DIP is a pair of literals. We use “Point” however to match the terminology of “Unique Implication Points”. More precisely, a DIP is defined to be a pair $\{x,y\}$ of literals such that all paths in the conflict graph to the vertex $\perp$ pass through at least one of $x$ and $y$ and such that neither $x$ nor $y$ is a UIP. In contrast, UIP is a single literal that separates the decision literal from the conflict.
We use the example in Figure 1 to illustrate the concept of DIPs and their potential applications. Recall that $x_{5}$ is the first UIP. In our applications, we are seeking DIPs between the first UIP and the conflict node $\perp$. An obvious DIP is the pair $\overline{x_{10}}$ and $x_{11}$, since it is immediately clear that any path from $x_{5}$ (or from $x_{1}$) must pass through one of $\overline{x_{10}}$ or $x_{11}$. On the other hand, the pair $\overline{x_{10}}$ and $x_{8}$ is not a DIP since there are paths from $x_{1}$ to $\perp$ that avoid these two literals; namely, any path that includes the edge from $\overline{x_{9}}$ to $x_{11}$. There are several other DIPs in Figure 1: a complete list is given in Figure 2.
Dual
Implication Extension Post-DIP Pre-DIP
Point (DIP) Variable Learned Clause Learned Clause
$\overline{x_{12}},x_{13}$ $z\leftrightarrow(\overline{x_{12}}\land x_{13})$$\lnot z$$\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor z$
$x_{11},x_{13}$ $z\leftrightarrow(x_{11}\land x_{13})$$\lnot z$$\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor z$
$\overline{x_{10}},x_{11}$ $z\leftrightarrow(\overline{x_{10}}\land x_{11})$$\lnot z$$\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor z$
$\overline{x_{9}},x_{11}$ $z\leftrightarrow(\overline{x_{9}}\land x_{11})$$\lnot z\lor\lnot y_{4}$$\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor z$
$x_{8},\overline{x_{9}}$ $z\leftrightarrow(x_{8}\land\overline{x_{9}})$$\lnot z\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}$$\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor z$
Figure 2: The complete list of DIPs for the conflict graph of Figure 1. The DIP-learnable clauses involve the new extension variable $z$ that can be introduced for that DIP. (The extension clauses defining $z$ must also be learned; e.g., for the first line, the extension clauses express that $z\leftrightarrow\overline{x_{12}}\land x_{13}$.)
Figure 2 also shows how a DIP pair can be used to introduce a variable $z$ via extension, and the associated clauses that can be learned. For example, in the third line, the new extension variable $z$ is introduced with the three clauses $\lnot z\lor\lnot x_{10}$, $\lnot z\lor x_{11}$ and $x_{10}\lor\lnot x_{11}\lor\lnot z$ which express the condition $z\leftrightarrow(\lnot x_{10}\land x_{11})$. From the conflict graph, this allows inferring the clauses $\lnot z$ (the “post-DIP” learned clause) and $\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor z$ (the “pre-DIP” learned clause). Since the post-DIP learned clause does not have any variables from lower decision levels, we can also infer the 2-clause $x_{10}\lor\lnot x_{11}$ either instead of or in addition to introducing $z$ and the pre- and post-DIP clauses. Introducing 2-clauses in this way might be helpful for CDCL solvers that do special processing of 2-clauses; for instance, in the work of Bacchus [5] or the recent work of Biere et al. [9] or Buss et al. [12].
In general, for any literals $a,b$ that form a DIP in the above fashion, we may introduce an extension variable $z\leftrightarrow a\land b$, and learn (1) a pre-DIP clause of the form $\lnot f\lor\lnot C\lor z$ (i.e. $f\land C\rightarrow z$), where $f$ is the first UIP and $C$ the set of literals from previous decision levels that have an edge in the conflict graph to any literal appearing after the first UIP and no later than the DIP pair; and (2) a post-DIP clause of the form $\lnot z\lor\lnot D$ (i.e. $z\land D\rightarrow\perp$), where $D$ contains those literals from previous decision levels with an edge to any literal appearing strictly after the DIP pair.
For example, the last line of the table in Figure 2 shows the case $z\leftrightarrow(x_{8}\land\overline{x_{9}})$, where $f$ is $x_{5}$, and $C$ and $D$ are $\lnot y_{1}\land y_{3}\land y_{4}$ and $y_{4}\land y_{5}\land\lnot y_{6}$, respectively. In Section 5.2 we discuss many possible ways that DIP extension variables and clauses may be introduced into CDCL solvers. The rest of this section describes how to find DIPs.
4.1 An Algorithm for Finding DIPs
A conventional CDCL algorithm maintains a trail of the literals set true, in the order they were set, and this allows finding the UIP very quickly. Finding the DIPs is much more complex than finding the UIP, and requires several traversals of the conflict graph between the first-UIP and the conflict node. Nonetheless, it is possible to find all DIPs very quickly, even in linear time.
We recast the DIP-finding problem in terms of a general directed graph $G$. Let $G=(V,E)$ be a directed graph with two distinguished vertices $s$ and $t$. We assume that $s$ is 2-connected to $t$ in that there is a pair of vertex-disjoint paths $\pi_{1}$ and $\pi_{2}$ from $s$ to $t$ that share no vertices apart from $s$ and $t$. By the vertex-version of Menger’s theorem [28], either there are in fact three vertex-disjoint paths from $s$ to $t$ or there is at least one pair of vertices $\{a,b\}$ with neither $a$ nor $b$ in $\{s,t\}$ so that every path from $s$ to $t$ passes through at least one of $a$ or $b$. Such a pair $\{a,b\}$, if it exists, is called a Two Vertex Bottleneck (TVB).
Our goal is to find all possible TVBs efficiently and in linear time. An algorithm for this is discussed in our companion paper [11]; for space reasons, we give only an abbreviated discussion of the algorithm here. The first step is to find two vertex disjoint paths from $s$ to $t$: this is done by greedily finding a path from $s$ to $t$ and then using an augmenting path construction to find two vertex-disjoint paths from $s$ to $t$. An example is shown in Figure 3, where the two paths are $a_{0},\ldots,a_{\ell}$ and $b_{0},\ldots,b_{k}$ where $a_{0}=b_{0}=s$ and $a_{\ell}=b_{k}=t$. These paths are called $\pi_{a}$ and $\pi_{b}$, respectively. Henceforth, a path is a directed path without any repeated nodes. The internal vertices of a path $\pi$ are the vertices on $\pi$ other than the first and last vertices. Two paths are said to be vertex-disjoint if they have no internal vertices in common. A path $\pi$ avoids $\pi_{a}$ and $\pi_{b}$ if it is vertex-disjoint from both paths.
$s$$a_{1}{=}a^{\prime}_{1}$$a_{2}{=}a^{\prime}_{2}$$a_{3}{=}a^{\prime}_{3}$$a_{4}$$a_{5}{=}a^{\prime}_{4}$$b_{1}{=}b^{\prime}_{1}$$b_{2}{=}b^{\prime}_{2}$$b_{3}$$b_{4}$$b_{5}{=}b^{\prime}_{3}$$t$$c$$d$
Figure 3: Nodes $a_{4}$, $b_{3}$ and $b_{4}$ are bypassed and thus not in any TVB. The path $b_{2},c,d,a_{4}$ is a crossing separator that prevents any of $a_{1}$, $a_{2}$ and $a_{3}$ from being paired with $b_{5}$ to form a TVB pair. All other pairs $\{a^{\prime}_{i},b^{\prime}_{j}\}$ are valid TVB pairs.
Once the two vertex-disjoint paths are fixed, we can state the following definitions and theorems:
Definition 4.1.
[11] A node $a_{i}$ on $\pi_{a}$ is bypassed if there are $j<i<j^{\prime}$ and a path $\pi$ from $a_{j}$ to $a_{j^{\prime}}$ such that $\pi$ avoids $\pi_{a}$ and $\pi_{b}$. A node $b_{i}$ being bypassed is defined similarly.
Definition 4.2.
Two nodes $a_{i}$ and $b_{j}$ have a crossing separator if there are nodes $a_{i^{\prime}}$ and $b_{j^{\prime}}$ joined by a path $\pi$ that avoids both $\pi_{a}$ and $\pi_{b}$ such that either (a) $i^{\prime}<i$ and $j^{\prime}>j$ and $\pi$ is a path from $a_{i^{\prime}}$ to $b_{j^{\prime}}$, or (b) $i^{\prime}>i$ and $j^{\prime}<j$ and $\pi$ is a path from $b_{j^{\prime}}$ to $a_{i^{\prime}}$.
Theorem 4.3.
[11] For $0<i<\ell$ and $0<j<k$, the two nodes $a_{i}$ and $b_{j}$ form a two-vertex bottleneck (TVB) if and only if $a_{i}$ and $b_{j}$ do not have a crossing separator and neither $a_{i}$ nor $b_{j}$ is bypassed.
Theorem 4.3 is proved in [11]. The theorem holds for both acyclic and cyclic directed graphs; however, for our applications to CDCL we are interested only in acyclic graphs since the conflict graph is always acyclic.
A consequence of Theorem 4.3 is that the set of all TVBs can be compactly represented in linear size, even though there can be quadratically many TVBs. Let $a^{\prime}_{1},\ldots a^{\prime}_{\ell^{\prime}}$ be the subsequence of the internal nodes $a_{1}.\ldots,a_{\ell-1}$ of path $\pi_{a}$ that, according to the conditions of Theorem 4.3, are in at least one TVB pair. Let $b^{\prime}_{1},\ldots,b^{\prime}_{k^{\prime}}$ be the corresponding subsequence of the internal nodes of $\pi_{b}$. Then, for each $a^{\prime}_{i}$ there are $m\leq n$ such that $a^{\prime}_{i}$ forms a TVB pair with each $b^{\prime}_{j}$ with $m\leq j\leq n$. Dually, for each $b^{\prime}_{i}$ there are $m\leq n$ such that $b^{\prime}_{i}$ forms a TVB pair with each $a^{\prime}_{j}$ with $m\leq j\leq n$.
Example 4.4.
In the conflict graph of Figure 1, consider the portion of the graph between the first UIP $x_{5}$ and the contradiction $\perp$. We can take path $\pi_{a}$ to be $x_{5},x_{7},x_{8},x_{11},\overline{x_{12}},\perp$ and path $\pi_{b}$ to be $x_{5},\overline{x_{9}},\overline{x_{10}},x_{13},\perp$. The node $x_{7}$ is bypassed by the path $x_{5},\overline{x_{6}},x_{8}$, so it cannot be part of a TVB. There are two crossing separator paths: the first is the edge from $\overline{x_{9}}$ to $x_{11}$; the second is the edge from $x_{11}$ to $x_{13}$. Therefore, the possible TVB pairs can be described in a table as:
Node forms a TVB Node forms a TVB
on $\pi_{a}$ pair with on $\pi_{b}$ pair with
$x_{8}$ $\overline{x_{9}}$ $\overline{x_{9}}$ $x_{8},x_{11}$
$x_{11}$ $\overline{x_{9}},\overline{x}_{{10}},x_{13}$ $\overline{x_{10}}$ $x_{11}$
$\overline{x_{12}}$ $x_{13}$ $x_{13}$ $x_{11},\overline{x}_{12}$
In this example, one node, $x_{7}$, was bypassed. It is also possible that non-bypassed nodes are eliminated just by the crossing separators. For example, if there were an additional edge from $x_{7}$ to $\overline{x_{10}}$, then the crossing separator condition would imply that $x_{8}$ and $x_{11}$ are not part of any TVB pair. In this case, $x_{8}$ and $x_{11}$ would not be included among the $a^{\prime}_{i}$ nodes.
Theorem 4.3 is used in [11] to give an efficient, linear time algorithm for finding all TVBs. The algorithm has five phases. The first two phases find two vertex-disjoint paths from $s$ to $t$; the next phase scans the graph from $t$ to $s$ to discover all relevant paths that avoid $\pi_{a}$ and $\pi_{b}$; the fourth phase uses this to discard bypassed nodes and collect information on crossing separators; finally the fifth phase computes the compressed representation of all possible TVBs. Full details are given [11], which is available online in preprint form. In our experiments with the implication graphs constructed by the underlying CDCL solver, the time overhead in finding the TVBs is negligible.
5 Extension Variables from Dual Implication Points
This section discusses how DIPs can be used for introducing extension variables and implement the ERCL method for learning ER clauses in CDCL solver. We first present an example.
5.1 An example with a grid Tseitin principle
Given a graph where every vertex has a charge, a number which is 00 or $1$, a Tseitin formula [37] is created by considering one variable per edge, and adding one constraint per vertex $v$ expressing that the sum of the variables of all edges incident to $v$ modulo 2 is equal to its charge. The CNF version that we consider converts each xor constraint into clauses by simple enumeration of falsifying assignments. It is easy to see that the formula is unsatisfiable if and only if the sum of all charges is even. Here we consider as a graph the $3\times 3$-grid depicted in Figure 4 and only one vertex has charge 1, so the clauses are unsatisfiable.
[…]
$e_{1}\oplus e_{3}$$=$ 1 $e_{5}\oplus e_{7}\oplus e_{10}$$=$ 0
$e_{1}\oplus e_{2}\oplus e_{4}$$=$ 0 $e_{8}\oplus e_{11}$$=$ 0
$e_{2}\oplus e_{5}$$=$ 0 $e_{9}\oplus e_{11}\oplus e_{12}$$=$ 0
$e_{3}\oplus e_{6}\oplus e_{8}$$=$ 0 $e_{10}\oplus e_{12}$$=$ 0
$e_{4}\oplus e_{6}\oplus e_{7}\oplus e_{9}$$=$ 0
Figure 4: An instance of the $3\times 3$-grid Tseitin principle. The nodes are assigned a polarity in $\{0,1\}$. The edges are labeled with variables.
The first steps of the CDCL solver are as follows: First $e_{1}$ is set true as a decision literal, and $\overline{e_{3}}$ is (unit) propagated. Second, $e_{2}$ is set true as a decision literal, and $\overline{e_{4}}$ and $e_{5}$ are propagated. Third, $e_{6}$ is set true as a decision literal, and $e_{8}$ and $e_{11}$ are propagated. Fourth, $e_{7}$ is set true as a decision literal, and $\overline{e_{9}}$, $\overline{e_{10}}$, and $e_{12}$ are propagated. This gives a contradiction, since the clause $e_{10}\lor\overline{e_{12}}$ (one of the two clauses from $e_{10}\oplus e_{12}=0$) is falsified. Figure 5(a) shows the complete conflict graph at this point.
Examining the conflict graph at decision level 4, the first UIP is $e_{7}$ and there are two DIPs available, $\{\overline{e_{9}},\overline{e_{10}}\}$ and $\{\overline{e_{10}},e_{12}\}$. Selecting the former DIP, we introduce a new variable $x$ by extension as $x\leftrightarrow\overline{e_{9}}\land\overline{e_{10}}$. We can learn the additional pre- and post-DIP clauses: $e_{4}\lor\overline{e_{5}}\lor\overline{e_{6}}\lor\overline{e_{7}}\lor x$ and $\overline{x}\lor\overline{e_{11}}$.
We next backtrack to decision level 3, unsetting $e_{7}$, $e_{9}$, $e_{10}$ and $e_{12}$. Unit propagation at decision level 3 sets the new literal $x$ and the first UIP $\overline{e_{7}}$ false and then sets literals $e_{9}$, $e_{10}$, and $\overline{e_{12}}$ true. This yields a contradiction with the clause $\overline{e_{10}}\lor e_{12}$. In the conflict graph at decision level 3, the first UIP is $e_{6}$ and there are two DIPs available, namely $\{\overline{e_{7}},\overline{e_{12}}\}$ and $\{e_{10},\overline{e_{12}}\}$. If we select the first one, then we introduce a new literal $y$ defined by extension as $y\leftrightarrow\overline{e_{7}}\land\overline{e_{12}}$ and can in addition learn the clauses $e_{3}\lor e_{4}\lor\overline{e_{5}}\lor\overline{e_{6}}\lor y$ and $\overline{e_{5}}\lor\overline{y}$.
We do not carry this example further, but note that our experiments show that Tseitin tautologies (not just on grid graphs) are examples where our experiments show the DIP clause learning method is especially effective.
$e_{1}$evel 1:$\overline{e_{3}}$$e_{2}$evel 2:$\overline{e_{4}}$$e_{5}$$e_{6}$evel 3:$e_{8}$$e_{11}$$e_{7}$evel 4:$\overline{e_{9}}$$e_{12}$$\overline{e_{10}}$$\perp$(a):
$e_{1}$evel 1:$\overline{e_{3}}$$e_{2}$evel 2:$\overline{e_{4}}$$e_{5}$$e_{6}$evel 3:$e_{8}$$e_{11}$$\overline{x}$$\overline{e_{7}}$$e_{10}$$e_{9}$$\overline{e_{12}}$$\perp$(b):
Figure 5: The complete conflict graph from the example; (a) at the first conflict and (b) at the second conflict. Boxed literals are decision literals. Implications at the top decision level are in bold.
It is interesting to relate the Figure 5(b) to Theorem 4.3 on DIPs. In this example, there is only one way to choose the two vertex-disjoint paths. Namely, to let $\pi_{a}$ and $\pi_{b}$ be the paths $e_{6},e_{8},e_{11},\overline{x},\overline{e_{7}},e_{10},\perp$ and $e_{6},e_{9},\overline{e_{12}},\perp$. The edge from $e_{6}$ to $\overline{e_{7}}$ bypasses $e_{8}$, $e_{11}$ and $\overline{x}$; and the edge from $e_{11}$ to $\overline{e_{12}}$ is a crossing separator. (The edge from $\overline{e_{7}}$ to $e_{9}$ is a “vacuous” crossing separator that does not actually remove any possible DIP pairs.)
It should be evident from this example that many conflicts have DIPs; indeed our experiments reported below show that approximately 2/3 of the conflicts have at least one DIP and, very frequently there are quite a few choices for DIPs.
5.2 Extending CDCL with Dual Implication Points
The use of DIPs in conflict analysis opens a large spectrum of possibilities. This section discusses some of them, with particular attention to the techniques that we have implemented. Our present implementation, xMapleLCM, is a flexible framework that allows one to implement extended-resolution based techniques in a simple way. It offers a set of clearly-specified functions that facilitate determining which extension variables to add, performing the corresponding addition, the possible posterior deletion or replacing definitions in clauses. Additionally, a set of heuristic choices to control when and how these steps are performed are provided, and replacing them by custom ones is a smooth task. We want to remark that there are many more possible strategies for using DIPs than could be discussed. We believe that in this paper we merely scratched the surface of heuristics for exploiting DIPs, which is an indication of the potential of this approach.
Choice of DIP. As mentioned, there is possibly a quadratic number of DIPs. Even though we could learn multiple DIPs at every conflict, with their corresponding lemmas, we decided to choose only one. The first possibility we considered is to learn the DIP that is closest to the conflict, as we do with UIPs. This may often create a short post-DIP conflict but a long pre-DIP clause. Therefore, we considered the possibility of choosing a DIP that splits the conflict graph into two balanced regions. Ideally, that would result in two equally long pre and post-DIP clauses. Finally, we also implemented choosing a random DIP, to check whether any of the other two schemes could outperform a random strategy. These heuristics are referred to as closest, middle, and random, respectively.
Filtering out bad-quality DIPs. Learning a DIP whenever we find one would be too prolific and overwhelm the solver. In order to determine whether the DIP chosen in the previous step had to be discarded, the first possibility we explored was again inspired by 1-UIP learning, where learning glue clauses, i.e. having LBD equal to $2$, is a desired situation. The first filtering mechanism we implemented discarded all DIPs that did not have a glue post-DIP clause. Another possibility we considered is to wait for a DIP to occur a certain number of times before using it in DIP-based learning. In our implementation, we tried $20$ and $5$ occurrences, as the minimum threshold to introduce a DIP. A third possibility is to use the activity-based heuristic of the literals in the DIP to assess its quality: DIPs whose literals have high decision-heuristic scores should be prioritized. In our implementation, we check whether the activity level of the current DIP is higher than the average activity level of the 20 most recently encountered DIPs; if so, the current DIP is a candidate for DIP-learning, otherwise it is discarded. We refer to these various techniques as glue, occ5 and act, respectively.
Learning pre-DIP and post-DIP clauses. In our implementation, we only considered two variants: one that always learns both the pre- and the post-DIP clauses (2-clause) and one that only learns the post-DIP clause (1-clause).
Backjumping and asserting clauses. Recall that when a new DIP extension variable $z$ is introduced, it is possible to learn a pre-DIP clause of the form $\lnot f\lor\lnot C\lor z$ and a post-DIP clause of the form $\lnot z\lor\lnot D$, where $f$ is the first UIP and where $C$ and $D$ are conjunctions of literals that were set true at lower levels. (Note that $C$ and $D$ may have literals in common.) Letting $\ell_{C}$ and $\ell_{D}$ be the maximum of the levels at which literals in $C$ and $D$ (respectively) were set, our implementation always backtracks to level $\ell_{D}$. This makes the post-DIP clause asserting, so $\lnot z$ is set at decision level $\ell_{D}$. Furthermore, if $\ell_{C}\leq\ell_{D}$ and the pre-DIP clause is learned, then it is asserting and $\lnot f$ is set by unit propagation at decision level $\ell_{D}$, as it should be.
It would make no sense to backtrack to level $\ell_{C}$ when $\ell_{C}<\ell_{D}$ since then neither $\lnot z$ nor $\lnot f$ would be propagated. However, another possible strategy would be to backtrack to the maximum of the decision levels $\ell_{C}$ and $\ell_{D}$. This would mean $\lnot z$ and $\lnot f$ are both propagated. The disadvantage of this when $\ell_{C}>\ell_{D}$ is that it would mean $\lnot z$ is asserted by the post-DIP clause at level $\ell_{C}$, whereas it could have been propagated at the previous decision level $\ell_{D}$. This breaks a usual invariant for CDCL solvers. This would only be possible in a solver that permits chronological backtracking [30, 29]; xMapleLCM, however, does not support this.
The previous reasoning needs some clarification for the case when the extension variable $z$ with definition $z\leftrightarrow l_{1}\land l_{2}$ we want to introduce already exists. If $z$ is undefined or defined at the current decision level nothing changes. We know it cannot be true at some previous level, because otherwise $l_{1}$ and $l_{2}$ would have been propagated at that level, and not in the current one as all literals belonging to a DIP do. If it is false at some previous level, then we have no guarantee that the pre or the post-DIP clause is asserting at any decision level. Fortunately, that rarely happens in practice. However, we can always perform standard 1UIP learning or try to apply DIP-based conflict analysis starting with the clause $z\lor\lnot l_{1}\lor\lnot l_{2}$ that we can guarantee is conflicting at the current decision level.
Replacing literals by extended variables. In xMapleLCM, every time a new lemma is learned, we try to replace some of its literals by extended variables. An extended variable $z\leftrightarrow l_{1}\lor l_{2}$ allows one to replace a lemma of the form $l_{1}\lor l_{2}\lor C$ by $z\lor C$. This is done by checking, for all pairs of literals in the lemma that appear in some extended variable definition whether they are part of the same definition. Lemmas that are too long or have large LBD are discarded to mitigate the cost of this operation.
If one wants to introduce literal $\lnot z$ and obtain some reduction in formula size, the lemma should be of the form $\lnot l_{1}\lor C$ and there should be another clause of the form $\lnot l_{2}\lor C$. In this case, they can be replaced by a single clause $\lnot z\lor C$. However, this situation can be expensive to detect since one has to traverse the whole database looking for a certain clause, and this operation should be repeated for every literal in the lemma. For this reason, xMapleLCM does not implement this.
Deleting extended variables. As it happens with lemma learning, where keeping too many lemmas slows down unit propagation, managing too many extended variables might also be counterproductive. Again following the analogy with learned lemmas, which are useful at some point of the search but might become inactive after a while, it is natural to think that extended variables follow the same behavior. All in all, it seems mandatory to consider the deletion of extended variables.
Deletion of variables in xMapleLCM is scheduled to be performed every $1000$ conflicts. At that point, several strategies are possible: delete all variables, delete the ones with a minimum decision-heuristic activity, or delete the worst $k\%$ variables according to some criterion (e.g. their decision-heuristic activity). In our implementation, we do the latter with $k=50$. Note that variables appearing in the right-hand side of an extended-variable definition cannot be deleted. This is addressed by maintaining a counter for every variable that corresponds to the number of definitions where it is involved.
6 Experimental Evaluation
We have implemented the DIP-based clause learning schemes described in Section 5.2 on top of the xMapleLCM ER framework [^3] All sources used for this evaluation can be found in https://github.com/chjon/xMapleSAT/tree/main.. We started our experimental evaluation running a variant of DIP-based learning on all benchmarks of the 2023 SAT Competition [6] and comparing it to MapleLCM [25], the CDCL SAT solver on which it is based. Even though there did not seem to be a systematic improvement on all benchmarks, a few families with important speedups were identified. We start this section by analyzing the impact of DIP-based learning on these families and then move to final considerations about the performance on the overall 2023 SAT competition benchmarks.
6.1 Performance Analysis Methodology
For each family, we start by describing the problem they encode. After that, we report on the performance of a variety of state-of-the-art solvers, each with some distinguished characteristic: Kissat 3.1.1 [8] (an extremely efficient CDCL solver), CryptoMiniSat 5.11 [36] (support for XOR reasoning), SBVA+CaDiCaL [14] (introduction of new variables via SBVA and winner of the main track of the 2023 SAT Competition), GlucoseER [3] (extended-resolution based CDCL solver) and a variant of xMapleLCM-DIP, which we refer to as the baseline, that we describe next.
Finally, we evaluate the impact of the different techniques explained in Section 5.2. In particular, we first consider as a baseline a version that (i) finds DIPs in the middle of the conflict graph, (ii) only adds a DIP if it has occurred at least 20 times and (iii) always learns the both pre- and post-DIP learning clauses. Different variants can be obtained by changing only one of the three previous design decisions at a time. Regarding the type of DIP used, we analyze the performance of closest and random, that is, the systems whose only difference w.r.t. baseline is that the type of DIP is changed. Regarding the criterion used to discard a DIP, we analyze the glue, act and occ5 configurations, where the latter discards any DIP that has not occurred at least 5 times. Finally, the 1-clause variant only learns the post-DIP clause. Our goal is to understand why certain configurations do not work well on particular families.
6.1.1 Matching of Properly Intersecting Intervals
We are given a sequence of numbers $(a_{1},\ldots,a_{n})$, initially all set to zero, and a set of operations, each consisting of assigning a certain number to a contiguous subsequence of the $a_{i}$’s, defined by an interval. The goal is to perform each operation exactly once while maximizing the number of pairs of consecutive numbers $(a_{i},a_{i+1})$ that are different at the end of the process. Given some additional conditions, this can be formulated as maximum bipartite matching problem on a certain graph. By removing some edges from this graph, an unsatisfiable problem is generated. A more detailed description can be found in [6].
We downloaded the 23 instances submitted to the 2023 SAT Competition, using the Global Benchmark Database [20], and observed that no system without extended-resolution reasoning could solve any benchmark in a time limit of 1 hour. On the other hand, there are 14 benchmarks that GlucoseER could solve, of which our DIP-based variants could solve 7. Detailed results for our variants can be seen in Table 1.
Baseline 1-Clause Closest Rand Glue Occ5 Act
PROPERLY INTERSECTING INTERVALS FORMULAS
SOLVED 6 5 7 7 0 0 0
BEST 1 1 2 3 0 0 0
Median (secs.) 2777 2566 2358 2189 3600 3600 3600
TSEITIN FORMULAS
SOLVED 11 13 13 11 0 15 3
BEST 2 5 0 0 0 8 0
Median (secs.) 897 628 2707 2782 3600 425 3600
(X)ORIFIED RANDOM $k$-xor FORMULAS
SOLVED 10 8 11 9 0 9 7
BEST 5 2 1 2 0 1 0
Median (secs.) 430 803 1080 820 3600 936 3112
Table 1: Performance of DIP-learning variants on three sets of formulas. Median running times are in seconds and, for each row, the best performing system is in bold.
For this particular type of formulas, the rand variant performs the best, but detailed results indicate that we cannot claim it is systematically faster. One clear conclusion we can infer is that glue, occ5 and act exhibit very poor performance. A more detailed analysis on the runs of these systems reveal that, at least in these benchmarks, both act and occ5 apply DIP learning in many more conflicts than the other variants, whereas the distinctive characteristic of glue is that the percentage of decisions on extended variables was extremely low. As can be seen at the end of this section, the latter is a situation than tends to indicate the DIP learning is not helping the solver in this context.
6.1.2 Tseitin Formulas
These formulas have already been described in Section 5.1. In this section, we consider unsatisfiable formulas generated by CNFgen [24]. We started generating instances with grids of size $n\times n$ and ran all systems with a time limit of 300 seconds in order to analyze how different solvers scale. The performance of all systems can be seen on the left plot in the first row of Figure 6 (note the logarithmic scale on the $y$ axis). A point $(x,y)$ in a solver line indicates that the solver took $y$ seconds to solve the Tseitin formula on grid size $x\times x$.
Figure 6: Performance analysis of different solvers on Tseitin formulas.
These formulas are easy for CryptoMiniSat, since the application of Gaussian elimination in a preprocessing step solves them. Regarding the rest of the systems, only our DIP-based tool and GlucoseER have good scaling properties. We want to remark that this behavior is not unique to Tseitin formulas on the grid. We generated Tseitin formulas over random $4$-regular and $6$-regular graphs and the conclusions are the same as can be seen in the other two plots in the first row of Figure 6. These results start suggesting that, even though GlucoseER and our DIP-based learning are quite different approaches to integrating ER into CDCL solvers, they seem to perform well on the same sets of benchmarks.
In order to understand whether the behavior of xMapleLCM-DIP was polynomial, we generated more challenging instances. Results can be seen in the left plot on the second row of Figure 6, with logarithmic scale on the $y$ axis. There is little doubt that the runtime of our solver ends up being exponential w.r.t. the size of the problem. However, we went one step further and studied how large the generated DRAT proofs were. In the same row, but on the right, we show the number of resolution steps in the proof. Despite having no theoretical support for that, it seems that although the solver takes exponential time in finding a proof, its size might be polynomial w.r.t. the problem size. That would indicate that our problem is having search heuristics that are not good enough to quickly find a short proof.
Finally, as we did in the previous section, we want to study the effect of the possible variants of DIP-based learning. For that purpose, we selected 5 challenging benchmarks of each type (grid, 4-regular, 6-regular) and evaluated the 7 DIP variants that we want to examine. Results can again be found in Table 1.
For Tseitin formulas, occ5 performs much better than the rest. Its only difference is that the minimum number of occurrences for a DIP to be introduced is smaller; this suggests that, in these instances, being more aggressive in the introduction of DIP extension variables provides a competitive advantage. On the other hand, glue solves no instance. We realized that DIP-learning was hardly ever applied, that is, requiring glue clauses was too restrictive here. Regarding act, whose performance is similarly poor, the main noticeable difference in its behavior is exhibiting a very large number of conflicts where DIP learning is applied, between 20 and 50%, which is around ten times more than the variants that performed best.
6.1.3 (X)Orified Random k-XOR Formulas
One last family where DIP-based systems perform very well are random $k$-xor formulas, where orification or xorification [7] has been applied, i.e., replacing variables by (x)ors of fresh variables. We used CNFgen to obtain 4 sets of formulas: xor constraints of length 3 applying orification with 2 and 3 variables, and also with xorification of 2 and 3 variables. All formulas we consider are unsatisfiable, the number of variables before (x)orification is equal to the number of clauses, and we increase this number to get progressively more difficult formulas.
We studied the scaling properties of all systems by running them on increasingly larger benchmarks with a time limit of 300 seconds. Results can be seen in Figure 7, where the $y$ axis has logarithmic scale. Maybe surprisingly, CryptoMiniSat can only benefit from its preprocessing step in one of the families (2-xorified). Also, SBVA-CaDiCaL performs very well in two of them (2 and 3-orified). Again, both GlucoseER and our DIP-based implementations show the best average overall performance over all competitors.
Figure 7: Performance analysis of different solvers on (x)orified random $k$-xor formulas.
Finally, an analysis of the impact of the different DIP-learning variants can be found in Table 1. Our baseline implementation seems to perform the best, but there are several other configurations that perform very well. On the negative side, glue did not solve any such formula within the time limit. The identifying characteristic of that version is that the percentage of conflicts with DIP-learning is very low compared to the others (around 0.1%).
6.2 Performance Analysis on 2023 SAT Competition Benchmarks
Our experimental evaluation concludes with the lessons we have learned from executing our DIP-variants [^4] We do not present a comparison with state-of-the-art solvers like Kissat or CaDiCaL since there are not many lessons to be learned. In a few words, our MapleLCM baseline CDCL solver is slower that those systems, and adding DIP-reasoning to it does not bridge this gap. on all benchmarks from the 2023 SAT Competition. We first report on the overhead caused by the DIP detection algorithm and the subsequent additional work to retrieve the clauses to be learned. For our baseline DIP-based system, where two clauses are learned and hence is the most computationally demanding method, in 4.5% of the benchmarks the DIP-related work represented between 10 and 15% of the total runtime; in 11.5% of the benchmarks between 5 and 10%; in 35.5% between 2.5 and in 48.5% less than 2.5%. These data show that DIP computation does not significantly slow down the solver. Another interesting information concerns the percentage of conflicts where there is at least one DIP, which was on average 63%. This implies that we do need to have filtering mechanisms to discard some of them. Otherwise, the search would be totally dominated by DIPs.
A cactus plot with the results of executing all our DIP variants on the 2023 SAT Competition benchmarks with a time limit of one hour is in the left plot of Figure 8. Three variants are much worse than the rest: act, occ5 and glue, whereas the other are relatively similar. These three configurations have a much larger percentage of conflicts where DIP-based learning is applied than the rest, being around 30% in act and around 5% in the other two. In the best performing variant this is around 1%, which seems to be a very low number and one might wonder whether this is enough to have any effect at all. This can be answered by analyzing the percentage of decisions on extended variables. For all benchmarks where DIP-based systems outperform the baseline CDCL solver, this number is surprisingly large, being usually over 10% of the decisions, whereas for most of the benchmarks is very low (in 70% of the benchmarks less than 1% of decisions are on extended variables). Moreover, this does not only happen in problems with a very low number of initial variables, where the introduction of a few extended variables could quickly dominate the decision heuristic.
This is remarkable since it shows that already existing decision heuristics like VSIDS or LRB somehow infer whether the newly introduced variables improve the behavior of the system. This is why we implemented, on top of our baseline DIP system, a procedure that computes the percentage of decisions on extended variables. If after a given number of conflicts it is still lower than 3%, it discontinues DIP learning and performs 1-UIP learning from that moment onwards. This is a very preliminary step in the direction of trying to automate the decision of whether to apply DIP reasoning or not, but the outcome, which is found in the right plot of Figure 8, is very positive.
Figure 8: Performance of DIP variants on 2023 SAT Competition benchmarks.
7 Conclusions and Future Work
We introduce a novel extended resolution clause learning (ERCL) algorithm, that when implemented on top of a CDCL solver, turns out to be beneficial for a variety of problems, in particular Tseitin, random k-xor and interval matching formulas. This is remarkable since automating powerful proof systems is well-known to be a difficult task. Somewhat surprisingly, the only previously existing attempt to incorporate ER into CDCL performs similarly on the same instances. Considering the different nature of the two methods, this clearly deserves further study. Further, we also introduce a new heuristic that allows our ERCL solver xMapleSAT to perform similarly to the baseline CDCL solver MapleLCM, thus being able to get the best of both worlds, i.e., the benefit of ERCL, without sacrificing performance of the CDCL solver on, say, SAT Competition Main Track 2023 instances. As future work, we plan to investigate a variety of machine learning based heuristics (e.g., branching) specialized for the ERCL method. Also, the use of $k$-IPs with $k>2$ is part of our next steps. On the theoretical side, challenging questions like determining whether DIP- or $k$-IP based ER simulates unrestricted ER are going to be central to our research efforts.

### v2

@@
 Extended Resolution Clause Learning via Dual Implication Points
-Abstract
+Abstract.
 We present a new extended resolution clause learning (ERCL) algorithm, implemented as part of a conflict-driven clause-learning (CDCL) SAT solver, wherein new variables are dynamically introduced as definitions for Dual Implication Points (DIPs) in the implication graph constructed by the solver at runtime.
 DIPs are generalizations of unique implication points and can be informally viewed as a pair of dominator nodes, from the decision variable at the highest decision level to the conflict node, in an implication graph.
-We perform extensive experimental evaluation to establish the efficacy of our ERCL method, implemented as part of the MapleLCM SAT solver and dubbed xMapleLCM, against several leading solvers including the baseline MapleLCM, as well as CDCL solvers such as Kissat 3.1.1, CryptoMiniSat 5.11, and SBVA+CaDiCaL, the winner of SAT Competition 2023.
-We show that xMapleLCM outperforms these solvers on Tseitin and XORified formulas.
-We further compare xMapleLCM with GlucoseER, a system that implements extended resolution in a different way, and provide a detailed comparative analysis of their performance.
-keywords:
+We perform extensive experimental evaluation to establish the efficacy of our ERCL method, implemented as part of the MapleLCM SAT solver and dubbed xMapleLCM, against several leading solvers such as Kissat, CryptoMiniSat, and SBVA-CaDiCaL, the winner of SAT Competition 2023.
+We show that xMapleLCM outperforms these solvers on Tseitin, XORified formulas and Interval Matching problems.
+We further compare xMapleLCM with GlucosER, a system that implements extended resolution in a different way, and provide a detailed comparative analysis of their performance.
+Key words and phrases:
 SAT, Extended resolution, Conflict Analysis
-category:
-[…]
-1 Introduction
-Over the last several years, Conflict-Driven Clause-Learning (CDCL) SAT solvers have had a dramatic impact on many fields including software engineering [23], security [39], and AI [22, 21].
+Albert Oliveras is supported by grant PID2021-122830OB-C43, funded by MCIN/AEI/ 10.13039/501100011033 and by “ERDF: A way of making Europe” and by Barcelogic through research grant C-11423.
+1. Introduction
+Over the last several years, Conflict-Driven Clause-Learning (CDCL) SAT solvers have had a dramatic impact on many fields including software engineering [Kro21], security [XA05], and AI [KS92, Kau06].
 As solvers continue to be adopted in increasing complex settings, the demand for greater efficiency and reasoning power by users continues unabated.
-While developers continue to improve CDCL SAT solvers, it is simultaneously true that these solvers are provably no more powerful than the relatively weak general resolution (Res) proof system [1, 32], and therefore are fundamentally limited.
+While developers continue to improve CDCL SAT solvers, it is simultaneously true that the base solvers are provably no more powerful than the relatively weak general resolution (Res) proof system [AFT11a, PD10], and therefore are fundamentally limited.
 Hence, solver developers have been actively researching novel algorithms that implement stronger proof systems that go beyond Res.
-Examples of such algorithms include satisfaction-driven clause-learning (SDCL) SAT solvers [17, 31], bounded variable addition (BVA) [26], symmetry breaking [35], and extended resolution (ER) solvers such as GlucoseER [3].
+Examples of such algorithms include satisfaction-driven clause-learning (SDCL) SAT solvers [HKB19, OLW+23], bounded variable addition (BVA) [MHB12], symmetry breaking [Sak21], and extended resolution (ER) solvers such as GlucosER [AKS10].
 Continuing this trend of strong proof system implementations, we present a new extended resolution clause learning (ERCL) algorithm, incorporated into a CDCL SAT solver, where new variables are dynamically introduced as definitions for dual implication points (DIPs) in the implication graph constructed by the solver at run time.
@@
 Finally, we need heuristics that specialize the above mentioned methods in a variety of ways, such as clause learning and clause deletion policies that are based on different kinds of DIPs.
-We implement all of these methods as part of the MapleLCM solver [25], and refer to the resulting solver as xMapleLCM.
+We implement all of these methods as part of the MapleLCM solver [LLX+17], and refer to the resulting solver as xMapleLCM.
 In fact our proposed ERCL method, and its implementation, is very general and easily extensible thus encouraging future exploration and specialization efforts with a variety of heuristics.
 Contributions.
-  1. 1.
+  1. (1)
     DIP: We introduce the concept of dual implication points (DIP), a generalization of UIPs in conflict graphs.
 We also came up with an algorithm that computes them in linear time.
-However, due to space limitations and to the non-trivial nature of the procedure, we discuss this in a separate paper [11].
-  2. 2.
+However, due to space limitations and to the non-trivial nature of the procedure, we discuss this in a separate paper [BGO24].
+  2. (2)
     ERCL Algorithm: We introduce a highly parameterizable DIP-based ERCL algorithm.
@@
 This flexibility is crucial in adapting the procedure to different scenarios, unlike previous methods that couple CDCL with extended resolution.
-  3. 3.
+  3. (3)
     xMapleLCM: We present a highly extensible and general ER framework as part of xMapleLCM, which allows developers to easily add their own new variable addition, ER clause learning/deletion, and branching policies.
 Given that DIP-based ERCL is highly flexible by nature, such a framework is necessary to quickly prototype new procedures.
-  4. 4.
-    Experiments: We perform extensive empirical evaluation and ablation studies on four different classes of instances, namely, SAT Competition 2023 Main Track, random-k-XOR, Tseitin, and interval matching, and compare xMapleLCM against leading solvers such as Kissat 3.1.1 [8], CryptoMiniSat 5.11 [36], SBVA CaDiCaL [14], and the extended-resolution solver GlucoseER [4].
-Results show that on the last three sets of hard combinatorial formulas, CDCL SAT solvers perform very poorly, whereas both xMapleLCM and GlucoseER excel on them.
+  4. (4)
+    Experiments: We perform extensive empirical evaluation and ablation studies on four different classes of instances, namely, SAT Competition 2023 Main Track, random $k$-xor, Tseitin, and Interval Matching, and compare xMapleLCM against leading solvers such as Kissat [BFFH20], CryptoMiniSat [SNC09], SBVA-CaDiCaL [HGH23], and the extended-resolution solver GlucosER [AS09].
+Results show that on the last three sets of hard combinatorial formulas, CDCL SAT solvers perform very poorly, whereas both xMapleLCM and GlucosER excel on them.
 Hence, we can now finally state that extended resolution can be added to CDCL and improve the performance of these solvers, and that there are at least two completely different ways to achieve this.
@@
 This technique enables xMapleLCM to perform similarly to MapleLCM on the SAT Competition 2023 Main Track instances.
-2 Related Work
+2. Related Work
 The idea of using ER in SAT solving has been studied in various forms in the literature for nearly two decades.
-The closest approach to ours is GlucoseER [3], where extended variables are introduced dynamically during the CDCL search: whenever two consecutive learned lemmas are of the form $\lnot l_{1}\lor C$ and $\lnot l_{2}\lor C$, with $l_{1}$ and $l_{2}$ being their UIPs, then an extended variable $z\leftrightarrow l_{1}\lor l_{2}$ is generated and any future lemma of the form $l_{1}\lor l_{2}\lor D$ is replaced by $z\lor D$.
-Our method differs significantly from GlucoseER in the way extended variables are identified: we choose DIPs, which can be seen as pairs of variables for which adding a definition would create a better first UIP (often written as 1UIP) in the conflict graph, whereas GlucoseER definitions are constructed using already existing UIPs.
-Also, unlike in GlucoseER, our approach does not always learn the standard 1UIP clause, but multiple clauses might be learned that take into account the newly introduced variable.
+The closest approach to ours is GlucosER [AKS10], where extended variables are introduced dynamically during the CDCL search: whenever two consecutive learned lemmas are of the form $\lnot l_{1}\lor C$ and $\lnot l_{2}\lor C$, with $l_{1}$ and $l_{2}$ being their UIPs, then an extended variable $z\leftrightarrow l_{1}\lor l_{2}$ is generated and any future lemma of the form $l_{1}\lor l_{2}\lor D$ is replaced by $z\lor D$.
+Hence, information across different conflict analysis steps is used to detect which extended variables to introduce.
+Our method differs significantly from GlucosER in the way extended variables are identified: we choose DIPs, which can be seen as pairs of variables for which adding a definition would create a better first UIP (often written as 1UIP) in the conflict graph, whereas GlucosER definitions are constructed using already existing UIPs.
+Our method focuses on a single conflict analysis step to extract an extended resolution variable, whereas GlucosER uses two steps.
+Also, unlike in GlucosER, our approach does not always learn the standard 1UIP clause, but multiple clauses might be learned that take into account the newly introduced variable.
 On the other hand, there are similarities at the heart of both procedures: a certain restriction of ER is considered and introduced variables are used to shorten subsequently found clauses.
-More recent uses of ER in SAT solvers are Bounded Variable Addition (BVA) [26] and its structured version SBVA [14], that due to clever strategies are able to identify new extended variables whose introduction can reduce the number of clauses.
+More recent uses of ER in SAT solvers are Bounded Variable Addition (BVA) [MHB12] and its structured version SBVA [HGH23], that due to clever strategies are able to identify new extended variables whose introduction can reduce the number of clauses.
 Essentially, the ultimate goal of BVA techniques is to reduce the size of the formula by reducing the number of clauses at the cost of adding a new variable.
 One big difference with our work is that this process is done only in a preprocessing step.
-Finally, one additional direction that has been researched is the development of a BDD-based solver to generate ER proofs [10].
-Other approaches aiming at improving CDCL solvers by allowing them to use a more powerful proof system are related to the Propagation Redundancy notion [16, 18], either via preprocessing steps [34] or via the use of the SDCL algorithm [19, 31].
+Finally, one additional direction that has been researched is the development of a BDD-based solver to generate ER proofs [BH23].
+Other approaches aiming at improving CDCL solvers by allowing them to use a more powerful proof system are related to the Propagation Redundancy notion [HKB17, HKB20], either via preprocessing steps [RHB22] or via the use of the SDCL algorithm [HKSB17, OLW+23].
 While these methods implement proof systems that are stronger than Res, many of them (without new variable addition) are known to be weaker than ER.
-3 Preliminaries
-We assume that the reader is familiar with the satisfiability (SAT) problem and the CDCL algorithm, and we refer her to the Handbook of Satisfiability for an excellent overview of these topics [27].
+3. Preliminaries
+We assume that the reader is familiar with the satisfiability (SAT) problem and the CDCL algorithm, and we refer her to the Handbook of Satisfiability for an excellent overview of these topics [MLM21].
 Below we focus on conflict analysis, which is the most relevant ingredient from the CDCL algorithm for this paper.
 We do so by means of the following example.
-Example 3.1.
+{exa}
 Consider the following clauses
@@
 Literals with this property are called Unique Implication Points (UIPs), of which we only have $x_{1}$ and $x_{5}$.
-Since $x_{5}$ is the one closest to the conflict we call it First Unique Implication Point (1UIP) [40].
+Since $x_{5}$ is the one closest to the conflict we call it First Unique Implication Point (1UIP) [ZMMM01].
 It is easy to see that if we set $x_{5}$ and the $y$ literals that enter the cut delimited by the blue line, unit propagation derives the same conflict.
 Hence, since they cannot be simultaneously true, we can learn $\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor x_{5}$.
-The quality of a lemma can be assessed by its Literal Block Distance (LBD) [4]: the number of different decision levels of the literals in the lemma.
+The quality of a lemma can be assessed by its Literal Block Distance (LBD) [AS09]: the number of different decision levels of the literals in the lemma.
 The lower the LBD, the better the lemma.
@@
 $x_{1}$$x_{2}$$\overline{x_{3}}$$x_{4}$$\overline{y_{2}}$$\overline{y_{1}}$$x_{5}$$y_{3}$$\overline{x_{6}}$$x_{7}$$\overline{x_{9}}$$y_{5}$$x_{8}$$\overline{y_{6}}$$x_{11}$$\overline{x_{10}}$$y_{4}$$\overline{x_{12}}$$x_{13}$$\perp$
-Figure 1: Conflict graph associated with Example 3.1.
+Figure 1.
+Conflict graph associated with Example 3.
 If 1UIP learning is applied, we generate lemma $\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor x_{5}$.
@@
 Given two clauses $l\lor C$ and $\lnot l\lor D$, the resolution inference rule allows one to derive the logical consequence $C\lor D$.
-It is well known that the lemma derived in Example 3.1 can be obtained via a series of resolution steps that start with the conflicting clause, and resolve with reasons of literals in the reverse order in which they were added to the assignment.
-In fact, it has been proved [33, 2] that resolution and CDCL (with restarts) are polynomially equivalent, and hence classes of formulas that are hard for resolution; e.g., the pigeonhole principle - PHP [15], or Tseitin formulas [38]) are also hard for CDCL SAT solvers.
+It is well known that the lemma derived in Example 3 can be obtained via a series of resolution steps that start with the conflicting clause, and resolve with reasons of literals in the reverse order in which they were added to the assignment.
+In fact, it has been proved [PD11, AFT11b] that resolution and CDCL (with restarts) are polynomially equivalent, and hence classes of formulas that are hard for resolution; e.g., the pigeonhole principle - PHP [Hak85], or Tseitin formulas [Urq87]) are also hard for CDCL SAT solvers.
 Extended Resolution (ER).
-Given two literals $l_{1}$ and $l_{2}$, the extended resolution [37] rule allows us to introduce clauses representing the definition $z\leftrightarrow l_{1}\lor l_{2}$.
-ER can be substantially more powerful than resolution; for instance, it allows polynomial size proofs of PHP [13].
+Given two literals $l_{1}$ and $l_{2}$, the extended resolution [Tse83] rule allows us to introduce clauses representing the definition $z\leftrightarrow l_{1}\lor l_{2}$.
+ER can be substantially more powerful than resolution; for instance, it allows polynomial size proofs of PHP [Coo76].
 Incorporating ER to CDCL solvers could potentially enable them to solve such formulas in polynomial time.
@@
 This is precisely the aim of this paper, namely, incorporating a restricted version of ER into CDCL.
-4 Dual Implication Points
+4. Dual Implication Points
 As discussed in the previous section, UIPs are crucial for conflict clause learning in the CDCL algorithm.
@@
 $x_{8},\overline{x_{9}}$ $z\leftrightarrow(x_{8}\land\overline{x_{9}})$$\lnot z\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}$$\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor z$
-Figure 2: The complete list of DIPs for the conflict graph of Figure 1.
+Figure 2.
+The complete list of DIPs for the conflict graph of Figure 1.
 The DIP-learnable clauses involve the new extension variable $z$ that can be introduced for that DIP.
@@
 Since the post-DIP learned clause does not have any variables from lower decision levels, we can also infer the 2-clause $x_{10}\lor\lnot x_{11}$ either instead of or in addition to introducing $z$ and the pre- and post-DIP clauses.
-Introducing 2-clauses in this way might be helpful for CDCL solvers that do special processing of 2-clauses; for instance, in the work of Bacchus [5] or the recent work of Biere et al.
-[9] or Buss et al.
-[12].
+Introducing 2-clauses in this way might be helpful for CDCL solvers that do special processing of 2-clauses; for instance, in the work of Bacchus [Bac02] or the recent work of Biere et al.
+[BFW23] or Buss et al.
+[BKW24].
 In general, for any literals $a,b$ that form a DIP in the above fashion, we may introduce an extension variable $z\leftrightarrow a\land b$, and learn (1) a pre-DIP clause of the form $\lnot f\lor\lnot C\lor z$ (i.e.
@@
 $z\land D\rightarrow\perp$), where $D$ contains those literals from previous decision levels with an edge to any literal appearing strictly after the DIP pair.
+The correctness of adding these two clauses is guaranteed because they can be generated via a sequence of resolution steps on the formula and the clauses that define the extended variable $z$.
 For example, the last line of the table in Figure 2 shows the case $z\leftrightarrow(x_{8}\land\overline{x_{9}})$, where $f$ is $x_{5}$, and $C$ and $D$ are $\lnot y_{1}\land y_{3}\land y_{4}$ and $y_{4}\land y_{5}\land\lnot y_{6}$, respectively.
@@
 The rest of this section describes how to find DIPs.
-4.1 An Algorithm for Finding DIPs
+4.1. An Algorithm for Finding DIPs
 A conventional CDCL algorithm maintains a trail of the literals set true, in the order they were set, and this allows finding the UIP very quickly.
@@
 We assume that $s$ is 2-connected to $t$ in that there is a pair of vertex-disjoint paths $\pi_{1}$ and $\pi_{2}$ from $s$ to $t$ that share no vertices apart from $s$ and $t$.
-By the vertex-version of Menger’s theorem [28], either there are in fact three vertex-disjoint paths from $s$ to $t$ or there is at least one pair of vertices $\{a,b\}$ with neither $a$ nor $b$ in $\{s,t\}$ so that every path from $s$ to $t$ passes through at least one of $a$ or $b$.
-Such a pair $\{a,b\}$, if it exists, is called a Two Vertex Bottleneck (TVB).
-Our goal is to find all possible TVBs efficiently and in linear time.
-An algorithm for this is discussed in our companion paper [11]; for space reasons, we give only an abbreviated discussion of the algorithm here.
+By the vertex-version of Menger’s theorem [Men27], either there are in fact three vertex-disjoint paths from $s$ to $t$ or there is at least one pair of vertices $\{a,b\}$ with neither $a$ nor $b$ in $\{s,t\}$ so that every path from $s$ to $t$ passes through at least one of $a$ or $b$.
+Such a pair $\{a,b\}$, if it exists, is called a Two Vertex Dominator (TVD).
+Our goal is to find all possible TVDs efficiently and in linear time.
+An algorithm for this is discussed in our companion paper [BGO24]; for space reasons, we give only an abbreviated discussion of the algorithm here.
 The first step is to find two vertex disjoint paths from $s$ to $t$: this is done by greedily finding a path from $s$ to $t$ and then using an augmenting path construction to find two vertex-disjoint paths from $s$ to $t$.
@@
 $s$$a_{1}{=}a^{\prime}_{1}$$a_{2}{=}a^{\prime}_{2}$$a_{3}{=}a^{\prime}_{3}$$a_{4}$$a_{5}{=}a^{\prime}_{4}$$b_{1}{=}b^{\prime}_{1}$$b_{2}{=}b^{\prime}_{2}$$b_{3}$$b_{4}$$b_{5}{=}b^{\prime}_{3}$$t$$c$$d$
-Figure 3: Nodes $a_{4}$, $b_{3}$ and $b_{4}$ are bypassed and thus not in any TVB.
-The path $b_{2},c,d,a_{4}$ is a crossing separator that prevents any of $a_{1}$, $a_{2}$ and $a_{3}$ from being paired with $b_{5}$ to form a TVB pair.
-All other pairs $\{a^{\prime}_{i},b^{\prime}_{j}\}$ are valid TVB pairs.
+Figure 3.
+Nodes $a_{4}$, $b_{3}$ and $b_{4}$ are bypassed and thus not in any TVD.
+The path $b_{2},c,d,a_{4}$ is a crossing separator that prevents any of $a_{1}$, $a_{2}$ and $a_{3}$ from being paired with $b_{5}$ to form a TVD pair.
+All other pairs $\{a^{\prime}_{i},b^{\prime}_{j}\}$ are valid TVD pairs.
 Once the two vertex-disjoint paths are fixed, we can state the following definitions and theorems:
-Definition 4.1.
-[11] A node $a_{i}$ on $\pi_{a}$ is bypassed if there are $j<i<j^{\prime}$ and a path $\pi$ from $a_{j}$ to $a_{j^{\prime}}$ such that $\pi$ avoids $\pi_{a}$ and $\pi_{b}$.
+{defi}
+[BGO24] A node $a_{i}$ on $\pi_{a}$ is bypassed if there are $j<i<j^{\prime}$ and a path $\pi$ from $a_{j}$ to $a_{j^{\prime}}$ such that $\pi$ avoids $\pi_{a}$ and $\pi_{b}$.
 A node $b_{i}$ being bypassed is defined similarly.
-Definition 4.2.
+{defi}
 Two nodes $a_{i}$ and $b_{j}$ have a crossing separator if there are nodes $a_{i^{\prime}}$ and $b_{j^{\prime}}$ joined by a path $\pi$ that avoids both $\pi_{a}$ and $\pi_{b}$ such that either (a) $i^{\prime}<i$ and $j^{\prime}>j$ and $\pi$ is a path from $a_{i^{\prime}}$ to $b_{j^{\prime}}$, or (b) $i^{\prime}>i$ and $j^{\prime}<j$ and $\pi$ is a path from $b_{j^{\prime}}$ to $a_{i^{\prime}}$.
-Theorem 4.3.
-[11] For $0<i<\ell$ and $0<j<k$, the two nodes $a_{i}$ and $b_{j}$ form a two-vertex bottleneck (TVB) if and only if $a_{i}$ and $b_{j}$ do not have a crossing separator and neither $a_{i}$ nor $b_{j}$ is bypassed.
-Theorem 4.3 is proved in [11].
+Theorem 1.
+[BGO24] For $0<i<\ell$ and $0<j<k$, the two nodes $a_{i}$ and $b_{j}$ form a two-vertex dominator (TVD) if and only if $a_{i}$ and $b_{j}$ do not have a crossing separator and neither $a_{i}$ nor $b_{j}$ is bypassed.
+Theorem 1 is proved in [BGO24].
 The theorem holds for both acyclic and cyclic directed graphs; however, for our applications to CDCL we are interested only in acyclic graphs since the conflict graph is always acyclic.
-A consequence of Theorem 4.3 is that the set of all TVBs can be compactly represented in linear size, even though there can be quadratically many TVBs.
-Let $a^{\prime}_{1},\ldots a^{\prime}_{\ell^{\prime}}$ be the subsequence of the internal nodes $a_{1}.\ldots,a_{\ell-1}$ of path $\pi_{a}$ that, according to the conditions of Theorem 4.3, are in at least one TVB pair.
+A consequence of Theorem 1 is that the set of all TVDs can be compactly represented in linear size, even though there can be quadratically many TVDs.
+Let $a^{\prime}_{1},\ldots a^{\prime}_{\ell^{\prime}}$ be the subsequence of the internal nodes $a_{1}.\ldots,a_{\ell-1}$ of path $\pi_{a}$ that, according to the conditions of Theorem 1, are in at least one TVD pair.
 Let $b^{\prime}_{1},\ldots,b^{\prime}_{k^{\prime}}$ be the corresponding subsequence of the internal nodes of $\pi_{b}$.
-Then, for each $a^{\prime}_{i}$ there are $m\leq n$ such that $a^{\prime}_{i}$ forms a TVB pair with each $b^{\prime}_{j}$ with $m\leq j\leq n$.
-Dually, for each $b^{\prime}_{i}$ there are $m\leq n$ such that $b^{\prime}_{i}$ forms a TVB pair with each $a^{\prime}_{j}$ with $m\leq j\leq n$.
-Example 4.4.
+Then, for each $a^{\prime}_{i}$ there are $m\leq n$ such that $a^{\prime}_{i}$ forms a TVD pair with each $b^{\prime}_{j}$ with $m\leq j\leq n$.
+Dually, for each $b^{\prime}_{i}$ there are $m\leq n$ such that $b^{\prime}_{i}$ forms a TVD pair with each $a^{\prime}_{j}$ with $m\leq j\leq n$.
+{exa}
 In the conflict graph of Figure 1, consider the portion of the graph between the first UIP $x_{5}$ and the contradiction $\perp$.
 We can take path $\pi_{a}$ to be $x_{5},x_{7},x_{8},x_{11},\overline{x_{12}},\perp$ and path $\pi_{b}$ to be $x_{5},\overline{x_{9}},\overline{x_{10}},x_{13},\perp$.
-The node $x_{7}$ is bypassed by the path $x_{5},\overline{x_{6}},x_{8}$, so it cannot be part of a TVB.
+The node $x_{7}$ is bypassed by the path $x_{5},\overline{x_{6}},x_{8}$, so it cannot be part of a TVD.
 There are two crossing separator paths: the first is the edge from $\overline{x_{9}}$ to $x_{11}$; the second is the edge from $x_{11}$ to $x_{13}$.
-Therefore, the possible TVB pairs can be described in a table as:
-Node forms a TVB Node forms a TVB
+Therefore, the possible TVD pairs can be described in a table as:
+Node forms a TVD Node forms a TVD
 on $\pi_{a}$ pair with on $\pi_{b}$ pair with
@@
 It is also possible that non-bypassed nodes are eliminated just by the crossing separators.
-For example, if there were an additional edge from $x_{7}$ to $\overline{x_{10}}$, then the crossing separator condition would imply that $x_{8}$ and $x_{11}$ are not part of any TVB pair.
+For example, if there were an additional edge from $x_{7}$ to $\overline{x_{10}}$, then the crossing separator condition would imply that $x_{8}$ and $x_{11}$ are not part of any TVD pair.
 In this case, $x_{8}$ and $x_{11}$ would not be included among the $a^{\prime}_{i}$ nodes.
-Theorem 4.3 is used in [11] to give an efficient, linear time algorithm for finding all TVBs.
+Theorem 1 is used in [BGO24] to give an efficient, linear time algorithm for finding all TVDs.
 The algorithm has five phases.
-The first two phases find two vertex-disjoint paths from $s$ to $t$; the next phase scans the graph from $t$ to $s$ to discover all relevant paths that avoid $\pi_{a}$ and $\pi_{b}$; the fourth phase uses this to discard bypassed nodes and collect information on crossing separators; finally the fifth phase computes the compressed representation of all possible TVBs.
-Full details are given [11], which is available online in preprint form.
-In our experiments with the implication graphs constructed by the underlying CDCL solver, the time overhead in finding the TVBs is negligible.
-5 Extension Variables from Dual Implication Points
+The first two phases find two vertex-disjoint paths from $s$ to $t$; the next phase scans the graph from $t$ to $s$ to discover all relevant paths that avoid $\pi_{a}$ and $\pi_{b}$; the fourth phase uses this to discard bypassed nodes and collect information on crossing separators; finally the fifth phase computes the compressed representation of all possible TVDs.
+Full details are given [BGO24], which is available online in preprint form.
+In our experiments with the implication graphs constructed by the underlying CDCL solver, the time overhead in finding the TVDs is negligible.
+5. Extension Variables from Dual Implication Points
 This section discusses how DIPs can be used for introducing extension variables and implement the ERCL method for learning ER clauses in CDCL solver. We first present an example.
-5.1 An example with a grid Tseitin principle
-Given a graph where every vertex has a charge, a number which is 00 or $1$, a Tseitin formula [37] is created by considering one variable per edge, and adding one constraint per vertex $v$ expressing that the sum of the variables of all edges incident to $v$ modulo 2 is equal to its charge.
+5.1. An example with a grid Tseitin principle
+Given a graph where every vertex has a charge, a number which is 00 or $1$, a Tseitin formula [Tse83] is created by considering one variable per edge, and adding one constraint per vertex $v$ expressing that the sum of the variables of all edges incident to $v$ modulo 2 is equal to its charge.
 The CNF version that we consider converts each xor constraint into clauses by simple enumeration of falsifying assignments.
@@
 $e_{4}\oplus e_{6}\oplus e_{7}\oplus e_{9}$$=$ 0
-Figure 4: An instance of the $3\times 3$-grid Tseitin principle. The nodes are assigned a polarity in $\{0,1\}$. The edges are labeled with variables.
+Figure 4. An instance of the $3\times 3$-grid Tseitin principle. The nodes are assigned a polarity in $\{0,1\}$. The edges are labeled with variables.
 The first steps of the CDCL solver are as follows: First $e_{1}$ is set true as a decision literal, and $\overline{e_{3}}$ is (unit) propagated.
@@
 $e_{1}$evel 1:$\overline{e_{3}}$$e_{2}$evel 2:$\overline{e_{4}}$$e_{5}$$e_{6}$evel 3:$e_{8}$$e_{11}$$\overline{x}$$\overline{e_{7}}$$e_{10}$$e_{9}$$\overline{e_{12}}$$\perp$(b):
-Figure 5: The complete conflict graph from the example; (a) at the first conflict and (b) at the second conflict.
+Figure 5.
+The complete conflict graph from the example; (a) at the first conflict and (b) at the second conflict.
 Boxed literals are decision literals.
 Implications at the top decision level are in bold.
-It is interesting to relate the Figure 5(b) to Theorem 4.3 on DIPs.
+It is interesting to relate the Figure 5(b) to Theorem 1 on DIPs.
 In this example, there is only one way to choose the two vertex-disjoint paths.
@@
 It should be evident from this example that many conflicts have DIPs; indeed our experiments reported below show that approximately 2/3 of the conflicts have at least one DIP and, very frequently there are quite a few choices for DIPs.
-5.2 Extending CDCL with Dual Implication Points
+5.2. Extending CDCL with Dual Implication Points
 The use of DIPs in conflict analysis opens a large spectrum of possibilities.
@@
 Another possibility we considered is to wait for a DIP to occur a certain number of times before using it in DIP-based learning.
-In our implementation, we tried $20$ and $5$ occurrences, as the minimum threshold to introduce a DIP.
+In our implementation, we tried different numbers as the minimum threshold to introduce a DIP.
 A third possibility is to use the activity-based heuristic of the literals in the DIP to assess its quality: DIPs whose literals have high decision-heuristic scores should be prioritized.
 In our implementation, we check whether the activity level of the current DIP is higher than the average activity level of the 20 most recently encountered DIPs; if so, the current DIP is a candidate for DIP-learning, otherwise it is discarded.
-We refer to these various techniques as glue, occ5 and act, respectively.
+We refer to these various techniques as glue, occ and act, respectively.
 Learning pre-DIP and post-DIP clauses.
@@
 This breaks a usual invariant for CDCL solvers.
-This would only be possible in a solver that permits chronological backtracking [30, 29]; xMapleLCM, however, does not support this.
-The previous reasoning needs some clarification for the case when the extension variable $z$ with definition $z\leftrightarrow l_{1}\land l_{2}$ we want to introduce already exists.
+This would only be possible in a solver that permits chronological backtracking [NR18, MB19]; xMapleLCM, however, does not support this.
+The previous reasoning needs some clarification for the case when the extension variable $z$ with definition $z\leftrightarrow l_{1}\land l_{2}$ we want to use has already been introduced.
 If $z$ is undefined or defined at the current decision level nothing changes.
@@
 Again following the analogy with learned lemmas, which are useful at some point of the search but might become inactive after a while, it is natural to think that extended variables follow the same behavior.
-All in all, it seems mandatory to consider the deletion of extended variables.
+All in all, it seems mandatory to consider the deletion of extended variables, which amounts to deletion all lemmas where they appear.
 Deletion of variables in xMapleLCM is scheduled to be performed every $1000$ conflicts.
@@
 This is addressed by maintaining a counter for every variable that corresponds to the number of definitions where it is involved.
-6 Experimental Evaluation
+6. Experimental Evaluation
 We have implemented the DIP-based clause learning schemes described in Section 5.2 on top of the xMapleLCM ER framework [^3] All sources used for this evaluation can be found in https://github.com/chjon/xMapleSAT/tree/main..
-We started our experimental evaluation running a variant of DIP-based learning on all benchmarks of the 2023 SAT Competition [6] and comparing it to MapleLCM [25], the CDCL SAT solver on which it is based.
+We started our experimental evaluation running a variant of DIP-based learning on benchmarks of the SAT Competition [BHI+23] and comparing it to MapleLCM [LLX+17], the CDCL SAT solver on which it is based.
 Even though there did not seem to be a systematic improvement on all benchmarks, a few families with important speedups were identified.
-We start this section by analyzing the impact of DIP-based learning on these families and then move to final considerations about the performance on the overall 2023 SAT competition benchmarks.
-6.1 Performance Analysis Methodology
-For each family, we start by describing the problem they encode.
-After that, we report on the performance of a variety of state-of-the-art solvers, each with some distinguished characteristic: Kissat 3.1.1 [8] (an extremely efficient CDCL solver), CryptoMiniSat 5.11 [36] (support for XOR reasoning), SBVA+CaDiCaL [14] (introduction of new variables via SBVA and winner of the main track of the 2023 SAT Competition), GlucoseER [3] (extended-resolution based CDCL solver) and a variant of xMapleLCM-DIP, which we refer to as the baseline, that we describe next.
-Finally, we evaluate the impact of the different techniques explained in Section 5.2.
-In particular, we first consider as a baseline a version that (i) finds DIPs in the middle of the conflict graph, (ii) only adds a DIP if it has occurred at least 20 times and (iii) always learns the both pre- and post-DIP learning clauses.
-Different variants can be obtained by changing only one of the three previous design decisions at a time.
-Regarding the type of DIP used, we analyze the performance of closest and random, that is, the systems whose only difference w.r.t.
+We start this section by analyzing the impact of DIP-based learning on these families and then move to final considerations about the performance on the overall SAT competition benchmarks.
+For each benchmark family, we start by describing the problem they encode.
+After that, we evaluate the impact of the different techniques explained in Section 5.2.
+In particular, we first consider as a baseline a version that (i) finds DIPs in the middle of the conflict graph, (ii) only adds a DIP if it has occurred at least 20 times and (iii) always learns both the pre- and post-DIP learning clauses.
+Different variants are obtained by changing only one of the three previous design decisions at a time.
+Regarding the type of DIP used, we analyze the performance of closest, random, and heuristic , that is, the systems whose only difference w.r.t.
 baseline is that the type of DIP is changed.
-Regarding the criterion used to discard a DIP, we analyze the glue, act and occ5 configurations, where the latter discards any DIP that has not occurred at least 5 times.
-Finally, the 1-clause variant only learns the post-DIP clause.
-Our goal is to understand why certain configurations do not work well on particular families.
-6.1.1 Matching of Properly Intersecting Intervals
+Regarding the criterion used to discard a DIP, we implemented the glue and the act configurations, but due to their poor performance we do not include these results.
+We present instead results about for the occ5 and occ50 configurations, which discard any DIP that has not occurred at least 5 or 50 times, respectively.
+Finally, we evaluated the performance of the 1-clause variant only learns the post-DIP clause.
+After that, we report on the performance of a variety of state-of-the-art solvers, each with some distinguished characteristic: Kissat 4.0.2 [BFFH20] (an extremely efficient CDCL solver), CryptoMiniSat 5.12.1 [SNC09] (support for XOR reasoning), SBVA-CaDiCaL [HGH23] (introduction of new variables via SBVA and winner of the main track of the 2023 SAT Competition), GlucosER [AKS10] (extended-resolution based CDCL solver) and the best available configuration for xMapleLCM.
+6.1. Tseitin Formulas
+These formulas have already been described in Section 5.1.
+In this section, we consider three types of unsatisfiable formulas generated by CNFgen [LENV17]: (i) Tseitin formulas on the grid, where the graph is a rectangular grid and each vertex is connected to its 4 neighbors, except for the vertices on the boundaries of the grid which have fewer neighbors, (ii) 4-regular Tseitin formulas, where the graph is a $4$-regular random one, and (iii) 6-regular Tseitin formulas built on top of $6$-regular random graphs.
+The time limit was set to 600 seconds per instance.
+Figure 6. Performance of different xMapleLCM configurations and state-of-the-art solvers on Tseitin formulas.
+The first row of Figure 6 shows the effect of changing the type of DIP that is used on our baseline solver.
+A point $(x,y)$ indicates that this particular variant took $y$ seconds (note the logarithmic scale on the $y$ axis) to solve the corresponding Tseitin formula of size $x$.
+We observe that there is no big difference for regular Tseitin formulas.
+However, for formulas on the grid, the middle DIP is the one that scales best, being the only variant able to solve all instances within the time limit.
+In the second row of the same figure, we can see how learning both the pre- and the post-DIP clauses (baseline in the plots) is comparable to only learning post-DIP clause (1-clause).
+Finally, the third row shows that changing the minimum number of times a DIP has to occur before we introduce it has some impact.
+50 seems to be too restrictive, whereas 20 (baseline) is slightly better than 5, which might end up adding too many DIPs.
+The last row of Figure 6 compares the baseline configuration of xMapleLCM with state-of-the-art tools.
+The conclusion is clear: our DIP-based xMapleLCM solver is the only one that can process all instances within the time limit.
+GlucosER scales well on regular formulas, but not on formulas on the grid.
+Despite these formulas are in principle trivial for CryptoMiniSat, since the application of Gaussian elimination in a preprocessing step solves them, this does not seem to work for large formulas on the grid.
+Finally, Kissat and SBVA-CaDiCaL perform poorly, solving only very small instances.
+In order to understand whether the behavior of xMapleLCM was polynomial for these formulas, we generated more challenging Tseitin formulas on the grid.
+Results can be seen in the left plot of Figure 7, with logarithmic scale on the $y$ axis.
+There is little doubt that the runtime of our solver ends up being exponential w.r.t.
+the size of the problem.
+However, we went one step further and studied how large the generated DRAT proofs were.
+On the right plot, we show the number of resolution steps in the proof.
+Despite having no theoretical support for that, it seems that although the solver takes exponential time in finding a proof, its size might be polynomial w.r.t.
+the problem size.
+That would indicate that our problem is using search heuristics that are not good enough to quickly find a short proof.
+Figure 7.
+Performance analysis of the baseline DIP-based solver on Tseitin formulas on the grid.
+The plot on the left represents running time (with logarithmic scale on $y$), whereas the one on the right shows the number of resolution steps in the generated proof.
+6.2. Xorified Random k-XOR Formulas
+Another family where DIP-based systems perform very well are random $k$-xor formulas, where xorification [BS09] has been applied, i.e., replacing variables by xors of fresh variables.
+We used CNFgen to obtain 2 sets of formulas: xor constraints of length 3 applying xorification with 2 and 3 variables.
+All formulas we consider are unsatisfiable, the number of variables before xorification is equal to the number of clauses, and we increase this number to get progressively more difficult formulas.
+A time limit of 600 seconds was used.
+The performance of the different xMapleLCM configurations can be seen in Figure 8.
+In the first row, we notice that there is not much difference if we change the DIP choice.
+A similar phenomenon can be observed in the second row, where the use of one or two learned DIP clauses does not have big impact on the performance.
+In the third row, we can see that changing the minimum number of occurrences of DIPs does affect the speed of the system: choosing 20 seems to give more consistent behavior.
+Finally, in the last row of Figure 8 we notice that CryptoMiniSat can only benefit from its preprocessing step when xorification with 2 variables is applied.
+Our xMapleLCM solver is the one the performs best with GlucosER in second place.
+Finally, both SBVA-CaDiCaL and Kissat are only able to solve small instances.
+Figure 8. Performance of different xMapleLCM configurations and state-of-the-arts solvers on XORified rand-kxor formulas.
+6.3. Matching of Properly Intersecting Intervals
 We are given a sequence of numbers $(a_{1},\ldots,a_{n})$, initially all set to zero, and a set of operations, each consisting of assigning a certain number to a contiguous subsequence of the $a_{i}$’s, defined by an interval.
@@
 By removing some edges from this graph, an unsatisfiable problem is generated.
-A more detailed description can be found in [6].
-We downloaded the 23 instances submitted to the 2023 SAT Competition, using the Global Benchmark Database [20], and observed that no system without extended-resolution reasoning could solve any benchmark in a time limit of 1 hour.
-On the other hand, there are 14 benchmarks that GlucoseER could solve, of which our DIP-based variants could solve 7.
-Detailed results for our variants can be seen in Table 1.
-Baseline 1-Clause Closest Rand Glue Occ5 Act
-PROPERLY INTERSECTING INTERVALS FORMULAS
-SOLVED 6 5 7 7 0 0 0
-BEST 1 1 2 3 0 0 0
-Median (secs.) 2777 2566 2358 2189 3600 3600 3600
-TSEITIN FORMULAS
-SOLVED 11 13 13 11 0 15 3
-BEST 2 5 0 0 0 8 0
-Median (secs.) 897 628 2707 2782 3600 425 3600
-(X)ORIFIED RANDOM $k$-xor FORMULAS
-SOLVED 10 8 11 9 0 9 7
-BEST 5 2 1 2 0 1 0
-Median (secs.) 430 803 1080 820 3600 936 3112
-Table 1: Performance of DIP-learning variants on three sets of formulas. Median running times are in seconds and, for each row, the best performing system is in bold.
-For this particular type of formulas, the rand variant performs the best, but detailed results indicate that we cannot claim it is systematically faster.
-One clear conclusion we can infer is that glue, occ5 and act exhibit very poor performance.
-A more detailed analysis on the runs of these systems reveal that, at least in these benchmarks, both act and occ5 apply DIP learning in many more conflicts than the other variants, whereas the distinctive characteristic of glue is that the percentage of decisions on extended variables was extremely low.
-As can be seen at the end of this section, the latter is a situation than tends to indicate the DIP learning is not helping the solver in this context.
-6.1.2 Tseitin Formulas
-These formulas have already been described in Section 5.1.
-In this section, we consider unsatisfiable formulas generated by CNFgen [24].
-We started generating instances with grids of size $n\times n$ and ran all systems with a time limit of 300 seconds in order to analyze how different solvers scale.
-The performance of all systems can be seen on the left plot in the first row of Figure 6 (note the logarithmic scale on the $y$ axis).
-A point $(x,y)$ in a solver line indicates that the solver took $y$ seconds to solve the Tseitin formula on grid size $x\times x$.
-Figure 6: Performance analysis of different solvers on Tseitin formulas.
-These formulas are easy for CryptoMiniSat, since the application of Gaussian elimination in a preprocessing step solves them.
-Regarding the rest of the systems, only our DIP-based tool and GlucoseER have good scaling properties.
-We want to remark that this behavior is not unique to Tseitin formulas on the grid.
-We generated Tseitin formulas over random $4$-regular and $6$-regular graphs and the conclusions are the same as can be seen in the other two plots in the first row of Figure 6.
-These results start suggesting that, even though GlucoseER and our DIP-based learning are quite different approaches to integrating ER into CDCL solvers, they seem to perform well on the same sets of benchmarks.
-In order to understand whether the behavior of xMapleLCM-DIP was polynomial, we generated more challenging instances.
-Results can be seen in the left plot on the second row of Figure 6, with logarithmic scale on the $y$ axis.
-There is little doubt that the runtime of our solver ends up being exponential w.r.t.
-the size of the problem.
-However, we went one step further and studied how large the generated DRAT proofs were.
-In the same row, but on the right, we show the number of resolution steps in the proof.
-Despite having no theoretical support for that, it seems that although the solver takes exponential time in finding a proof, its size might be polynomial w.r.t.
-the problem size.
-That would indicate that our problem is having search heuristics that are not good enough to quickly find a short proof.
-Finally, as we did in the previous section, we want to study the effect of the possible variants of DIP-based learning.
-For that purpose, we selected 5 challenging benchmarks of each type (grid, 4-regular, 6-regular) and evaluated the 7 DIP variants that we want to examine.
-Results can again be found in Table 1.
-For Tseitin formulas, occ5 performs much better than the rest.
-Its only difference is that the minimum number of occurrences for a DIP to be introduced is smaller; this suggests that, in these instances, being more aggressive in the introduction of DIP extension variables provides a competitive advantage.
-On the other hand, glue solves no instance.
-We realized that DIP-learning was hardly ever applied, that is, requiring glue clauses was too restrictive here.
-Regarding act, whose performance is similarly poor, the main noticeable difference in its behavior is exhibiting a very large number of conflicts where DIP learning is applied, between 20 and 50%, which is around ten times more than the variants that performed best.
-6.1.3 (X)Orified Random k-XOR Formulas
-One last family where DIP-based systems perform very well are random $k$-xor formulas, where orification or xorification [7] has been applied, i.e., replacing variables by (x)ors of fresh variables.
-We used CNFgen to obtain 4 sets of formulas: xor constraints of length 3 applying orification with 2 and 3 variables, and also with xorification of 2 and 3 variables.
-All formulas we consider are unsatisfiable, the number of variables before (x)orification is equal to the number of clauses, and we increase this number to get progressively more difficult formulas.
-We studied the scaling properties of all systems by running them on increasingly larger benchmarks with a time limit of 300 seconds.
-Results can be seen in Figure 7, where the $y$ axis has logarithmic scale.
-Maybe surprisingly, CryptoMiniSat can only benefit from its preprocessing step in one of the families (2-xorified).
-Also, SBVA-CaDiCaL performs very well in two of them (2 and 3-orified).
-Again, both GlucoseER and our DIP-based implementations show the best average overall performance over all competitors.
-Figure 7: Performance analysis of different solvers on (x)orified random $k$-xor formulas.
-Finally, an analysis of the impact of the different DIP-learning variants can be found in Table 1.
-Our baseline implementation seems to perform the best, but there are several other configurations that perform very well.
-On the negative side, glue did not solve any such formula within the time limit.
-The identifying characteristic of that version is that the percentage of conflicts with DIP-learning is very low compared to the others (around 0.1%).
-6.2 Performance Analysis on 2023 SAT Competition Benchmarks
-Our experimental evaluation concludes with the lessons we have learned from executing our DIP-variants [^4] We do not present a comparison with state-of-the-art solvers like Kissat or CaDiCaL since there are not many lessons to be learned.
-In a few words, our MapleLCM baseline CDCL solver is slower that those systems, and adding DIP-reasoning to it does not bridge this gap.
-on all benchmarks from the 2023 SAT Competition.
+A more detailed description can be found in [BHI+23].
+We downloaded the 23 instances submitted to the SAT Competition, using the Global Benchmark Database [IS19], and observed that no system without extended-resolution reasoning could solve any benchmark in a time limit of 2 hours.
+We used a much larger timeout than with the two previous families because for this family we are not able to construct increasingly larger benchmarks and see how the solvers scale.
+Table 1 reports on our experimental evaluation.
+In the first block we can see that the random and heuristic DIP choices do significantly improve upon our baseline solver.
+The second block row shows how the behavior of the heuristic configuration can be further improved by increasing the threshold of minimum DIP occurrences to 50.
+On the contrary, a smaller threshold of 5 significantly degrades the performance, whereas only learning the post-DIP clause slightly slows it down.
+Finally, the third block compares our best xMapleLCM configuration with state-of-the-art solvers.
+It is remarkable that, apart from GlucosER, no system is able to process any instance within the given time limit.
+Hence, the use of extended resolution is crucial in this set of benchmarks.
+Our comparison shows that GlucosER and xMapleLCM solve the same number of instances, the former being only slightly faster.
+Baseline Random Closest Heuristic
+SOLVED 6 13 10 13
+Average (secs.) 6011 4337 4818 4220
+Heuristic 1-clause Occ-5 Occ-50
+SOLVED 13 11 6 14
+Average (secs.) 4220 4995 6057 3913
+Occ-50 GlucosER Kissat CryptoMiniSat SBVA-CaDiCaL
+SOLVED 14 14 0 0 0
+Average (secs.) 3913 3622 7200 7200 7200
+Table 1. Average running times in seconds of DIP-learning variants on properly intersecting intervals formulas. The best performing systems are in bold.
+6.4. Performance Analysis on SAT Competition Benchmarks
+Our experimental evaluation concludes with the lessons we have learned from executing our DIP implementation on the SAT competition benchmarks.
+We used benchmarks from the 2023 and 2024 editions.
+In the latter, the number of instances from the Tseitin, Xorified random $k$-xor and Interval families we have described is much lower than in the 2023 edition, and hence we present results from the 2023 competition in order to show the possible gains or losses of adding DIP-based reasoning to solvers.
 We first report on the overhead caused by the DIP detection algorithm and the subsequent additional work to retrieve the clauses to be learned.
@@
 Otherwise, the search would be totally dominated by DIPs.
-A cactus plot with the results of executing all our DIP variants on the 2023 SAT Competition benchmarks with a time limit of one hour is in the left plot of Figure 8.
-Three variants are much worse than the rest: act, occ5 and glue, whereas the other are relatively similar.
-These three configurations have a much larger percentage of conflicts where DIP-based learning is applied than the rest, being around 30% in act and around 5% in the other two.
-In the best performing variant this is around 1%, which seems to be a very low number and one might wonder whether this is enough to have any effect at all.
-This can be answered by analyzing the percentage of decisions on extended variables.
-For all benchmarks where DIP-based systems outperform the baseline CDCL solver, this number is surprisingly large, being usually over 10% of the decisions, whereas for most of the benchmarks is very low (in 70% of the benchmarks less than 1% of decisions are on extended variables).
-Moreover, this does not only happen in problems with a very low number of initial variables, where the introduction of a few extended variables could quickly dominate the decision heuristic.
+In the left cactus plot of Figure 9 we can see that our baseline DIP-based xMapleLCM solver outperforms the CDCL-based MapleLCM that it is based on.
+When we analyzed the concrete benchmarks where our DIP-based solver outperforms MapleLCM and other CDCL-based solvers like Kissat, we observed that for all of them the percentage of decisions on extended variables is surprisingly large (over 10% of the decisions), whereas for the remaining benchmarks it is very low (in 70% of the benchmarks less than 1% of decisions are on extended variables).
+Moreover, this happens no matter how many initial variables there are, even though in problems with few number of initial variables the introduction of a few extended variables could quickly dominate the decision heuristic.
 This is remarkable since it shows that already existing decision heuristics like VSIDS or LRB somehow infer whether the newly introduced variables improve the behavior of the system.
@@
 If after a given number of conflicts it is still lower than 3%, it discontinues DIP learning and performs 1-UIP learning from that moment onwards.
-This is a very preliminary step in the direction of trying to automate the decision of whether to apply DIP reasoning or not, but the outcome, which is found in the right plot of Figure 8, is very positive.
-Figure 8: Performance of DIP variants on 2023 SAT Competition benchmarks.
-7 Conclusions and Future Work
-We introduce a novel extended resolution clause learning (ERCL) algorithm, that when implemented on top of a CDCL solver, turns out to be beneficial for a variety of problems, in particular Tseitin, random k-xor and interval matching formulas.
+This is a very preliminary step in the direction of trying to automate the decision of whether to apply DIP reasoning or not, but the outcome, which is found on the left cactus plot of Figure 9, is very positive, showing that a solver disabling DIP reasoning (xMapleLCM-disabling in the plot) when the percentage of decisions on extended variables is low is still able to solve Tseitin, random $k$-xor and Interval benchmarks, but the overall performance is not significantly worsened on the remaining instances.
+Finally, the right cactus plot shows the impact of changing the type of DIP into the disabling solver: the heuristic and middle DIPs, the latter being used in the baseline configuration, are much better than random and closest ones.
+Figure 9. Performance of DIP variants on 2023 SAT Competition benchmarks.
+Learned lessons from experimental evaluation. Let us finish this section by pointing out the main lessons we have learned from this experimental evaluation.
+  • We have identified three families of benchmarks for which the dynamic addition of extension variables improves upon CDCL in a systematic manner.
+This is an improvement w.r.t.
+the results in the GlucosER paper [AKS10], where a few initial benchmarks were identified but it was not proved that extended-resolution methods always outperform CDCL on increasingly more complex instances.
+  • We have proved that, even though the way extended variables are identified is completely different in xMapleLCM and GlucosER, they both behave well on the same instances.
+This is remarkable because our preliminary experiments revealed that randomly choosing arbitrary pairs of (non-DIP) literals to create definitions does not provide any speedup w.r.t.
+CDCL.
+Hence, both methods are able to identify semantic properties about concrete literals that make them suitable for creating an extended variable definition.
+  • We have learned that in most of the families where extended resolution works there is a large xor component.
+Additionally, when this is xor information is somehow hidden in the formula, our dynamic approach outperforms static methods like the ones in CryptoMiniSat and SBVA-CaDiCaL.
+  • Our ablation experiments have shown that the most important design choice is the one of determining which DIP is to be chosen.
+Among all the possibilities we have devised, middle and heuristic are the preferable ones.
+  • Finally, we have noticed that by inspecting the percentage of decisions on extended variables, one can detect whether our DIP-based is helping the solver.
+This has allowed us to further improve the behavior of our solver on sets of benchmarks where only a few of them are advantageous for extended resolution.
+7. Conclusions and Future Work
+We introduce a novel extended resolution clause learning (ERCL) algorithm that, when implemented on top of a CDCL solver, turns out to be beneficial for a variety of problems, in particular Tseitin, random $k$-xor and interval matching formulas.
 This is remarkable since automating powerful proof systems is well-known to be a difficult task.
-Somewhat surprisingly, the only previously existing attempt to incorporate ER into CDCL performs similarly on the same instances.
+Somewhat surprisingly, the only previously existing attempt to incorporate ER into CDCL performs reasonably well on the same instances.
+We have hence indentified classes of formulas for which extended resolution can be automated in at least two different ways and still outperform CDCL.
 Considering the different nature of the two methods, this clearly deserves further study.

### v3

@@
 SAT, Extended resolution, Conflict Analysis
-Albert Oliveras is supported by grant PID2021-122830OB-C43, funded by MCIN/AEI/ 10.13039/501100011033 and by “ERDF: A way of making Europe” and by Barcelogic through research grant C-11423.
+Albert Oliveras is supported by grant PID2021-122830OB-C43, funded by MCIN/AEI/ 10.13039/501100011033 and by “ERDF: A way of making Europe”; by grant PID2024-157044OB-C32, funded by MICIU /AEI /10.13039/501100011033 / FEDER, UE.
+and by Barcelogic through research grant C-11423.
 1. Introduction
-Over the last several years, Conflict-Driven Clause-Learning (CDCL) SAT solvers have had a dramatic impact on many fields including software engineering [Kro21], security [XA05], and AI [KS92, Kau06].
+Over the last several years, Conflict-Driven Clause-Learning (CDCL) SAT solvers have had a dramatic impact on many fields including software engineering [HandbookSATSoftwareVerification], security [xie2005security], and AI [satplan1, satplan2].
 As solvers continue to be adopted in increasing complex settings, the demand for greater efficiency and reasoning power by users continues unabated.
-While developers continue to improve CDCL SAT solvers, it is simultaneously true that the base solvers are provably no more powerful than the relatively weak general resolution (Res) proof system [AFT11a, PD10], and therefore are fundamentally limited.
+While developers continue to improve CDCL SAT solvers, it is simultaneously true that the base solvers are provably no more powerful than the relatively weak general resolution (Res) proof system [AFT11ClauseLearning, PipatsrisawatDarwiche:modernclauselearning], and therefore are fundamentally limited.
 Hence, solver developers have been actively researching novel algorithms that implement stronger proof systems that go beyond Res.
-Examples of such algorithms include satisfaction-driven clause-learning (SDCL) SAT solvers [HKB19, OLW+23], bounded variable addition (BVA) [MHB12], symmetry breaking [Sak21], and extended resolution (ER) solvers such as GlucosER [AKS10].
+Examples of such algorithms include satisfaction-driven clause-learning (SDCL) SAT solvers [HKB:RedundancySDCL, SDCLMaxSAT], bounded variable addition (BVA) [BVA], symmetry breaking [HandbookSATSymmetry], and extended resolution (ER) solvers such as GlucosER [Audemard2010ARO].
 Continuing this trend of strong proof system implementations, we present a new extended resolution clause learning (ERCL) algorithm, incorporated into a CDCL SAT solver, where new variables are dynamically introduced as definitions for dual implication points (DIPs) in the implication graph constructed by the solver at run time.
@@
 Finally, we need heuristics that specialize the above mentioned methods in a variety of ways, such as clause learning and clause deletion policies that are based on different kinds of DIPs.
-We implement all of these methods as part of the MapleLCM solver [LLX+17], and refer to the resulting solver as xMapleLCM.
+We implement all of these methods as part of the MapleLCM solver [MapleLCM], and refer to the resulting solver as xMapleLCM.
 In fact our proposed ERCL method, and its implementation, is very general and easily extensible thus encouraging future exploration and specialization efforts with a variety of heuristics.
@@
 We also came up with an algorithm that computes them in linear time.
-However, due to space limitations and to the non-trivial nature of the procedure, we discuss this in a separate paper [BGO24].
+However, due to space limitations and to the non-trivial nature of the procedure, we discuss this in a separate paper [BGO:TVBs].
   2. (2)
@@
   4. (4)
-    Experiments: We perform extensive empirical evaluation and ablation studies on four different classes of instances, namely, SAT Competition 2023 Main Track, random $k$-xor, Tseitin, and Interval Matching, and compare xMapleLCM against leading solvers such as Kissat [BFFH20], CryptoMiniSat [SNC09], SBVA-CaDiCaL [HGH23], and the extended-resolution solver GlucosER [AS09].
-Results show that on the last three sets of hard combinatorial formulas, CDCL SAT solvers perform very poorly, whereas both xMapleLCM and GlucosER excel on them.
-Hence, we can now finally state that extended resolution can be added to CDCL and improve the performance of these solvers, and that there are at least two completely different ways to achieve this.
-Moreover, a simple heuristic dynamically allows us to detect whether extended resolution is being helpful and turn back to standard CDCL in order to get the best of both worlds.
+    Experiments: We perform extensive empirical evaluation and ablation studies on four different classes of instances, namely, SAT Competition 2023 Main Track, random $k$-xor, Tseitin, and Interval Matching, and compare xMapleLCM against leading solvers such as Kissat [kissat], CryptoMiniSat [cryptominisat], SBVA-CaDiCaL [SBVA], and the extended-resolution solver GlucosER [LBDAudemardSimon].
+Results show that on the last three sets of hard combinatorial formulas, CDCL SAT solvers perform very poorly, whereas both xMapleLCM and GlucosER present much better performance.
+This demostrates that extended resolution can be added to CDCL and improve the performance of these solvers; GlucosER and our DIP-based methods give two different ways to achieve this.
+Moreover, for our DIP-based learning, a simple heuristic allows us to detect on-the-fly whether extended resolution is being helpful and revert to using standard CDCL in order to get the best of both worlds.
 This technique enables xMapleLCM to perform similarly to MapleLCM on the SAT Competition 2023 Main Track instances.
@@
 The idea of using ER in SAT solving has been studied in various forms in the literature for nearly two decades.
-The closest approach to ours is GlucosER [AKS10], where extended variables are introduced dynamically during the CDCL search: whenever two consecutive learned lemmas are of the form $\lnot l_{1}\lor C$ and $\lnot l_{2}\lor C$, with $l_{1}$ and $l_{2}$ being their UIPs, then an extended variable $z\leftrightarrow l_{1}\lor l_{2}$ is generated and any future lemma of the form $l_{1}\lor l_{2}\lor D$ is replaced by $z\lor D$.
+The closest approach to ours is GlucosER [Audemard2010ARO], where extended variables are introduced dynamically during the CDCL search: whenever two consecutive learned lemmas are of the form $\lnot l_{1}\lor C$ and $\lnot l_{2}\lor C$, with $l_{1}$ and $l_{2}$ being their UIPs, then an extended variable $z\leftrightarrow l_{1}\lor l_{2}$ is generated and any future lemma of the form $l_{1}\lor l_{2}\lor D$ is replaced by $z\lor D$.
 Hence, information across different conflict analysis steps is used to detect which extended variables to introduce.
@@
 On the other hand, there are similarities at the heart of both procedures: a certain restriction of ER is considered and introduced variables are used to shorten subsequently found clauses.
-More recent uses of ER in SAT solvers are Bounded Variable Addition (BVA) [MHB12] and its structured version SBVA [HGH23], that due to clever strategies are able to identify new extended variables whose introduction can reduce the number of clauses.
+More recent uses of ER in SAT solvers are Bounded Variable Addition (BVA) [BVA] and its structured version SBVA [SBVA], that due to clever strategies are able to identify new extended variables whose introduction can reduce the number of clauses.
 Essentially, the ultimate goal of BVA techniques is to reduce the size of the formula by reducing the number of clauses at the cost of adding a new variable.
 One big difference with our work is that this process is done only in a preprocessing step.
-Finally, one additional direction that has been researched is the development of a BDD-based solver to generate ER proofs [BH23].
-Other approaches aiming at improving CDCL solvers by allowing them to use a more powerful proof system are related to the Propagation Redundancy notion [HKB17, HKB20], either via preprocessing steps [RHB22] or via the use of the SDCL algorithm [HKSB17, OLW+23].
+Finally, one additional direction that has been researched is the development of a BDD-based solver to generate ER proofs [ExtendedResolutionBDDs].
+Other approaches aiming at improving CDCL solvers by allowing them to use a more powerful proof system are related to the Propagation Redundancy notion [ShortProofs, StrongExtensionFree], either via preprocessing steps [preprocessingPR] or via the use of the SDCL algorithm [PruningThroughSatisfaction, SDCLMaxSAT].
 While these methods implement proof systems that are stronger than Res, many of them (without new variable addition) are known to be weaker than ER.
+Another related work is [PD08Learning], where the authors introduce a new conflict analysis procedure that learns lemmas that contain two literals of the last decision level.
+These two literals are indeed a DIP.
+However, there can be many other candidate DIP’s that are not detected by the methods of [PD08Learning].
+Furthermore, as we will show in Section 4.1, their procedure might sometimes fail to detect any DIPs at all.
+An additional, important differences w.r.t.
+our work is that [PD08Learning] do not introduce a new variable representing the conjunction of the DIP members.
+Hence, this is not an extended resolution-based approach and, from the proof-complexity point of view, is not more powerful than standard CDCL or general resolution.
 3. Preliminaries
-We assume that the reader is familiar with the satisfiability (SAT) problem and the CDCL algorithm, and we refer her to the Handbook of Satisfiability for an excellent overview of these topics [MLM21].
+We assume that the reader is familiar with the satisfiability (SAT) problem and the CDCL algorithm, and we refer her to the Handbook of Satisfiability for an excellent overview of these topics [HandbookCDCL].
 Below we focus on conflict analysis, which is the most relevant ingredient from the CDCL algorithm for this paper.
@@
 Consider the following clauses
-$\begin{array}[]{llllll}(1)&y_{1}\lor\lnot x_{1}\lor x_{2}&(6)&\lnot x_{5}\lor x_{7}&(11)&\lnot x_{11}\lor x_{12}\\ (2)&\lnot x_{1}\lor\lnot x_{3}&(7)&x_{6}\lor\lnot x_{7}\lor x_{8}&(12)&x_{10}\lor\lnot x_{11}\lor x_{13}\\ (3)&y_{2}\lor\lnot x_{1}\lor x_{4}&(8)&\lnot y_{3}\lor\lnot y_{4}\lor\lnot x_{5}\lor\lnot x_{9}&(13)&x_{12}\lor\lnot x_{13}\\ (4)&\lnot y_{3}\lor\lnot x_{2}\lor x_{3}\lor\lnot x_{4}\lor x_{5}&(9)&\lnot y_{4}\lor x_{9}\lor\lnot x_{10}\\ (5)&y_{1}\lor\lnot x_{5}\lor\lnot x_{6}&(10)&\lnot y_{5}\lor y_{6}\lor\lnot x_{8}\lor x_{9}\lor x_{11}\\ \end{array}$
+[…]
 Assume that CDCL has constructed an assignment that contains, among others, literals $\{\lnot y_{1},\lnot y_{2},y_{3},y_{4},y_{5},\lnot y_{6}\}$.
@@
 A special conflict node $\perp$ with incoming edges $\{\overline{x_{12}},x_{13}\}$ represents conflicting clause (13).
-The graph clearly shows that if we set $x_{1}$ to true, together with the literals of previous decision levels (the $y$’s), we obtain a conflict.
-However, the same happens with $x_{5}$, since any path from $x_{1}$ to the conflict necessarily goes through $x_{5}$.
-Literals with this property are called Unique Implication Points (UIPs), of which we only have $x_{1}$ and $x_{5}$.
-Since $x_{5}$ is the one closest to the conflict we call it First Unique Implication Point (1UIP) [ZMMM01].
-It is easy to see that if we set $x_{5}$ and the $y$ literals that enter the cut delimited by the blue line, unit propagation derives the same conflict.
-Hence, since they cannot be simultaneously true, we can learn $\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor x_{5}$.
-The quality of a lemma can be assessed by its Literal Block Distance (LBD) [AS09]: the number of different decision levels of the literals in the lemma.
-The lower the LBD, the better the lemma.
-In our case, if we are at decision level 5, $y_{4}$ and $y_{6}$ belong to decision level 2, and $y_{5}$ to decision level 4, the LBD of the lemma is 3.
-$\Box$
 $x_{1}$$x_{2}$$\overline{x_{3}}$$x_{4}$$\overline{y_{2}}$$\overline{y_{1}}$$x_{5}$$y_{3}$$\overline{x_{6}}$$x_{7}$$\overline{x_{9}}$$y_{5}$$x_{8}$$\overline{y_{6}}$$x_{11}$$\overline{x_{10}}$$y_{4}$$\overline{x_{12}}$$x_{13}\perp$
@@
 White nodes belong to the current decision level, whereas blue ones are from previous decision levels.
+The graph clearly shows that if we set $x_{1}$ to true, together with the literals of previous decision levels (the $y$’s), we obtain a conflict.
+However, the same happens with $x_{5}$, since any path from $x_{1}$ to the conflict necessarily goes through $x_{5}$.
+Literals with this property are called Unique Implication Points (UIPs), of which we only have $x_{1}$ and $x_{5}$.
+Since $x_{5}$ is the one closest to the conflict we call it First Unique Implication Point (1UIP) [ConflictAnalysis].
+It is easy to see that if we set $x_{5}$ and the $y$ literals that enter the cut delimited by the blue line, unit propagation derives the same conflict.
+Hence, since they cannot be simultaneously true, we can learn $\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor x_{5}$.
+The quality of a lemma can be assessed by its Literal Block Distance (LBD) [LBDAudemardSimon]: the number of different decision levels of the literals in the lemma.
+The lower the LBD, the better the lemma.
+In our case, if we are at decision level 5, $y_{4}$ and $y_{6}$ belong to decision level 2, and $y_{5}$ to decision level 4, the LBD of the lemma is 3.
+$\Box$
 Resolution.
@@
 It is well known that the lemma derived in Example 3 can be obtained via a series of resolution steps that start with the conflicting clause, and resolve with reasons of literals in the reverse order in which they were added to the assignment.
-In fact, it has been proved [PD11, AFT11b] that resolution and CDCL (with restarts) are polynomially equivalent, and hence classes of formulas that are hard for resolution; e.g., the pigeonhole principle - PHP [Hak85], or Tseitin formulas [Urq87]) are also hard for CDCL SAT solvers.
+In fact, it has been proved [CDCLResolutionKnot, CDCLResolutionAtserias] that resolution and CDCL (with restarts) are polynomially equivalent, and hence classes of formulas that are hard for resolution; e.g., the pigeonhole principle - PHP [haken1985intractability], or Tseitin formulas [TseitinHardForResolution]) are also hard for CDCL SAT solvers.
 Extended Resolution (ER).
-Given two literals $l_{1}$ and $l_{2}$, the extended resolution [Tse83] rule allows us to introduce clauses representing the definition $z\leftrightarrow l_{1}\lor l_{2}$.
-ER can be substantially more powerful than resolution; for instance, it allows polynomial size proofs of PHP [Coo76].
+Given two literals $l_{1}$ and $l_{2}$, the extended resolution [tseitin1983complexity] rule allows us to introduce clauses representing the definition $z\leftrightarrow l_{1}\lor l_{2}$.
+ER can be substantially more powerful than resolution; for instance, it allows polynomial size proofs of PHP [PigeonholeExtendedResolution].
 Incorporating ER to CDCL solvers could potentially enable them to solve such formulas in polynomial time.
@@
 Point (DIP) Variable Learned Clause Learned Clause
-$\overline{x_{12}},x_{13}$ $z\leftrightarrow(\overline{x_{12}}\land x_{13})$$\lnot z$$\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor z$
-$x_{11},x_{13}$ $z\leftrightarrow(x_{11}\land x_{13})$$\lnot z$$\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor z$
-$\overline{x_{10}},x_{11}$ $z\leftrightarrow(\overline{x_{10}}\land x_{11})$$\lnot z$$\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor z$
+$\overline{x_{12}},x_{13}$ $\leftrightarrow(\overline{x_{12}}\land x_{13}) ¬z\lnot z$ $\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor z$
+$x_{11},x_{13}$ $\leftrightarrow(x_{11}\land x_{13}) ¬z\lnot z$ $\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor z$
+$\overline{x_{10}},x_{11}$ $\leftrightarrow(\overline{x_{10}}\land x_{11}) ¬z\lnot z$ $\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor z$
 $\overline{x_{9}},x_{11}$ $z\leftrightarrow(\overline{x_{9}}\land x_{11})$ $\lnot z\lor\lnot y_{4}$ $\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor z$
@@
 For example, in the third line, the new extension variable $z$ is introduced with the three clauses $\lnot z\lor\lnot x_{10}$, $\lnot z\lor x_{11}$ and $x_{10}\lor\lnot x_{11}\lor\lnot z$ which express the condition $z\leftrightarrow(\lnot x_{10}\land x_{11})$.
-From the conflict graph, this allows inferring the clauses $\lnot z$ (the “post-DIP” learned clause) and $\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor z$ (the “pre-DIP” learned clause).
+From the conflict graph, this allows inferring the clauses $\lnot z$ (the “post-DIP” learned clause) and $\lnot x_{5}\lor y_{1}\lor\lnot y_{3}\lor\lnot y_{4}\lor\lnot y_{5}\lor y_{6}\lor z$ (the “pre-DIP” learned clause), both of which will be formally defined in the next paragraph.
 Since the post-DIP learned clause does not have any variables from lower decision levels, we can also infer the 2-clause $x_{10}\lor\lnot x_{11}$ either instead of or in addition to introducing $z$ and the pre- and post-DIP clauses.
-Introducing 2-clauses in this way might be helpful for CDCL solvers that do special processing of 2-clauses; for instance, in the work of Bacchus [Bac02] or the recent work of Biere et al.
-[BFW23] or Buss et al.
-[BKW24].
+Introducing 2-clauses in this way might be helpful for CDCL solvers that do special processing of 2-clauses; for instance, in the work of Bacchus [Bacchus:EnhancingDP] or the recent work of Biere et al.
+[BFW:CadiBack] or Buss et al.
+[BKVW:DualDFS].
 In general, for any literals $a,b$ that form a DIP in the above fashion, we may introduce an extension variable $z\leftrightarrow a\land b$, and learn (1) a pre-DIP clause of the form $\lnot f\lor\lnot C\lor z$ (i.e.
@@
 The correctness of adding these two clauses is guaranteed because they can be generated via a sequence of resolution steps on the formula and the clauses that define the extended variable $z$.
+We can also prove their correctness with a more conflict-graph oriented argument: it should be obvious from the conflict graph that if we assert the 1UIP $f$ and all literals in $C$, we can obtain $a$ and $b$ via unit propagation.
+This means that $f\land C\rightarrow a\land b$ is a valid formula which, after replacing $a\land b$ by their definition $z$, becomes the pre-DIP clause $\lnot f\lor\lnot C\lor z$.
+Similarly, if we assert $a,b$ and all literals in $D$, we get a conflict by unit propagation.
+This means that $\lnot a\lor\lnot b\lor\lnot D$ is a valid clause.
+If we apply two resolution steps with the clauses $a\lor\lnot z$ and $b\lor\lnot z$ we obtain the post-DIP clause $\lnot z\lor\lnot D$.
 For example, the last line of the table in Figure 2 shows the case $z\leftrightarrow(x_{8}\land\overline{x_{9}})$, where $f$ is $x_{5}$, and $C$ and $D$ are $\lnot y_{1}\land y_{3}\land y_{4}$ and $y_{4}\land y_{5}\land\lnot y_{6}$, respectively.
 In Section 5.2 we discuss many possible ways that DIP extension variables and clauses may be introduced into CDCL solvers.
-The rest of this section describes how to find DIPs.
+The rest of this section describes how to find DIPs and its content is not strictly necessary to understand the rest of the paper.
 4.1. An Algorithm for Finding DIPs
+$x_{1}$$x_{3}$$x_{4}$$x_{2}$$x_{5}$$x_{6}$$x_{7}\perp$
+ Trail Derived clause $\perp$ $\lnot x_{5}\lor\lnot x_{6}\lor\lnot x_{7}$ $x_{5}$ $\lnot x_{3}\lor\lnot x_{4}\lor\lnot x_{6}\lor\lnot x_{7}$ $x_{7}$ $\lnot x_{3}\lor\lnot x_{4}\lor\lnot x_{6}\lor\lnot x_{2}$ $x_{6}$ $\lnot x_{3}\lor\lnot x_{4}\lor\lnot x_{2}$ $x_{4}$ $\lnot x_{3}\lor\lnot x_{1}\lor\lnot x_{2}$ $x_{3}$ $\lnot x_{1}\lor\lnot x_{2}$ $x_{2}$ $\lnot x_{1}$ $x_{1}$
+Figure 3.
+Conflict graph associated with Example 3, with the literals in blue being the only DIP.
+Next to it, the resulting trail and the sequence of intermediate clauses obtained if standard 1UIP learning is applied.
 A conventional CDCL algorithm maintains a trail of the literals set true, in the order they were set, and this allows finding the UIP very quickly.
+The idea is initialize a clause $C$ as the conflicting clause and, following the order of the trail from the most recent literal $l$ backwards, apply resolution between $C$ and the reason of $l$ to get a new clause $C$.
+This is repeated until $C$ only has only one literal of the current decision level.
+One could think that finding DIPs can be done by applying the same algorithm and stopping when $C$ has exactly two literals of the current decision level.
+This is what is done in [PD08Learning] but, as we now show, this procedure is not complete for detecting DIPs.
+{exa} Consider the set of clauses $\{\lnot x_{1}\lor x_{2},\allowbreak\;\lnot x_{1}\lor x_{3},\allowbreak\;\lnot x_{1}\lor x_{4},\allowbreak\;\lnot x_{3}\lor\lnot x_{4}\lor x_{5},\allowbreak\;\lnot x_{2}\lor x_{6},\allowbreak\;\lnot x_{2}\lor x_{7},\allowbreak\;\lnot x_{5}\lor\lnot x_{6}\lor\lnot x_{7}\}$.
+Setting decision $x_{1}$ and applying unit propagation finds a conflict.
+This is illustrated in the conflict graph of Figure 3, where the only DIP is the pair $\{x_{2},x_{5}\}$.
+In the same figure, we can see the corresponding trail and the intermediate clauses obtained in a standard conflict analysis procedure.
+Note that the procedure is unable to detect the DIP.
 Finding the DIPs is much more complex than finding the UIP, and requires several traversals of the conflict graph between the first-UIP and the conflict node.
@@
 We assume that $s$ is 2-connected to $t$ in that there is a pair of vertex-disjoint paths $\pi_{1}$ and $\pi_{2}$ from $s$ to $t$ that share no vertices apart from $s$ and $t$.
-By the vertex-version of Menger’s theorem [Men27], either there are in fact three vertex-disjoint paths from $s$ to $t$ or there is at least one pair of vertices $\{a,b\}$ with neither $a$ nor $b$ in $\{s,t\}$ so that every path from $s$ to $t$ passes through at least one of $a$ or $b$.
+By the vertex-version of Menger’s theorem [Menger:Kurventheorie], either there are in fact three vertex-disjoint paths from $s$ to $t$ or there is at least one pair of vertices $\{a,b\}$ with neither $a$ nor $b$ in $\{s,t\}$ so that every path from $s$ to $t$ passes through at least one of $a$ or $b$.
 Such a pair $\{a,b\}$, if it exists, is called a Two Vertex Dominator (TVD).
 Our goal is to find all possible TVDs efficiently and in linear time.
-An algorithm for this is discussed in our companion paper [BGO24]; for space reasons, we give only an abbreviated discussion of the algorithm here.
-The first step is to find two vertex disjoint paths from $s$ to $t$: this is done by greedily finding a path from $s$ to $t$ and then using an augmenting path construction to find two vertex-disjoint paths from $s$ to $t$.
+An algorithm for this is discussed in detail in our companion paper [BGO:TVBs]; for space reasons, we give only an abbreviated discussion of the algorithm here and direct the reader who wants a deep understanding of the algorithm to [BGO:TVBs].
+The first step is to find two vertex disjoint paths from $s$ to $t$: this is done by greedily finding a path from $s$ to $t$ and then using an augmenting path [^3] This can also be computed in linear time by applying Ford-Fulkerson’s algorithm to a small modification of the conflict graph in order look for a flow of 2.
+construction to find two vertex-disjoint paths from $s$ to $t$.
 An example is shown in Figure 4, where the two paths are $a_{0},\ldots,a_{\ell}$ and $b_{0},\ldots,b_{k}$ where $a_{0}=b_{0}=s$ and $a_{\ell}=b_{k}=t$.
@@
 A path $\pi$ avoids $\pi_{a}$ and $\pi_{b}$ if it is vertex-disjoint from both paths.
-$s$$a_{1}{=}a^{\prime}_{1}$$a_{2}{=}a^{\prime}_{2}$$a_{3}{=}a^{\prime}_{3}$$a_{4}$$a_{5}{=}a^{\prime}_{4}$$b_{1}{=}b^{\prime}_{1}$$b_{2}{=}b^{\prime}_{2}$$b_{3}$$b_{4}$$b_{5}{=}b^{\prime}_{3}$$t$$c$$d$
+[…]
 Figure 4.
@@
 {defi}
-[BGO24] A node $a_{i}$ on $\pi_{a}$ is bypassed if there are $j<i<j^{\prime}$ and a path $\pi$ from $a_{j}$ to $a_{j^{\prime}}$ such that $\pi$ avoids $\pi_{a}$ and $\pi_{b}$.
+[BGO:TVBs] A node $a_{i}$ on $\pi_{a}$ is bypassed if there are $j<i<j^{\prime}$ and a path $\pi$ from $a_{j}$ to $a_{j^{\prime}}$ such that $\pi$ avoids $\pi_{a}$ and $\pi_{b}$.
 A node $b_{i}$ being bypassed is defined similarly.
@@
 Theorem 1.
-[BGO24] For $0<i<\ell$ and $0<j<k$, the two nodes $a_{i}$ and $b_{j}$ form a two-vertex dominator (TVD) if and only if $a_{i}$ and $b_{j}$ do not have a crossing separator and neither $a_{i}$ nor $b_{j}$ is bypassed.
-Theorem 1 is proved in [BGO24].
+[BGO:TVBs] For $0<i<\ell$ and $0<j<k$, the two nodes $a_{i}$ and $b_{j}$ form a two-vertex dominator (TVD) if and only if $a_{i}$ and $b_{j}$ do not have a crossing separator and neither $a_{i}$ nor $b_{j}$ is bypassed.
+Theorem 1 is proved in [BGO:TVBs].
 The theorem holds for both acyclic and cyclic directed graphs; however, for our applications to CDCL we are interested only in acyclic graphs since the conflict graph is always acyclic.
@@
 In this case, $x_{8}$ and $x_{11}$ would not be included among the $a^{\prime}_{i}$ nodes.
-Theorem 1 is used in [BGO24] to give an efficient, linear time algorithm for finding all TVDs.
+Theorem 1 is used in [BGO:TVBs] to give an efficient, linear time algorithm for finding all TVDs.
 The algorithm has five phases.
 The first two phases find two vertex-disjoint paths from $s$ to $t$; the next phase scans the graph from $t$ to $s$ to discover all relevant paths that avoid $\pi_{a}$ and $\pi_{b}$; the fourth phase uses this to discard bypassed nodes and collect information on crossing separators; finally the fifth phase computes the compressed representation of all possible TVDs.
-Full details are given [BGO24], which is available online in preprint form.
+Full details are given [BGO:TVBs], which is available online in preprint form.
 In our experiments with the implication graphs constructed by the underlying CDCL solver, the time overhead in finding the TVDs is negligible.
@@
 5.1. An example with a grid Tseitin principle
-Given a graph where every vertex has a charge, a number which is 00 or $1$, a Tseitin formula [Tse83] is created by considering one variable per edge, and adding one constraint per vertex $v$ expressing that the sum of the variables of all edges incident to $v$ modulo 2 is equal to its charge.
+Given a graph where every vertex has a charge, a number which is 0 or $1$, a Tseitin formula [tseitin1983complexity] is created by considering one variable per edge, and adding one constraint per vertex $v$ expressing that the sum of the variables of all edges incident to $v$ modulo 2 is equal to its charge.
 The CNF version that we consider converts each xor constraint into clauses by simple enumeration of falsifying assignments.
-It is easy to see that the formula is unsatisfiable if and only if the sum of all charges is even.
+It is easy to see that the formula is unsatisfiable if and only if the sum of all charges is odd.
 Here we consider as a graph the $3\times 3$-grid depicted in Figure 5 and only one vertex has charge 1, so the clauses are unsatisfiable.
 […]
-$e_{1}\oplus e_{3}$$=$ 1 $e_{5}\oplus e_{7}\oplus e_{10}$$=$ 0
-$e_{1}\oplus e_{2}\oplus e_{4}$$=$ 0 $e_{8}\oplus e_{11}$$=$ 0
-$e_{2}\oplus e_{5}$$=$ 0 $e_{9}\oplus e_{11}\oplus e_{12}$$=$ 0
-$e_{3}\oplus e_{6}\oplus e_{8}$$=$ 0 $e_{10}\oplus e_{12}$$=$ 0
-$e_{4}\oplus e_{6}\oplus e_{7}\oplus e_{9}$$=$ 0
+$e_{1}\oplus e_{3}$== 1 $e_{5}\oplus e_{7}\oplus e_{10}$== 0
+$e_{1}\oplus e_{2}\oplus e_{4}$== 0 $e_{8}\oplus e_{11}$== 0
+$e_{2}\oplus e_{5}$== 0 $e_{9}\oplus e_{11}\oplus e_{12}$== 0
+$e_{3}\oplus e_{6}\oplus e_{8}$== 0 $e_{10}\oplus e_{12}$== 0
+$e_{4}\oplus e_{6}\oplus e_{7}\oplus e_{9}$== 0
 Figure 5. An instance of the $3\times 3$-grid Tseitin principle. The nodes are assigned a polarity in $\{0,1\}$. The edges are labeled with variables.
@@
 We do not carry this example further, but note that our experiments show that Tseitin tautologies (not just on grid graphs) are examples where our experiments show the DIP clause learning method is especially effective.
-$e_{1}$evel 1:$\overline{e_{3}}$$e_{2}$evel 2:$\overline{e_{4}}$$e_{5}$$e_{6}$evel 3:$e_{8}$$e_{11}$$e_{7}$evel 4:$\overline{e_{9}}$$e_{12}$$\overline{e_{10}}$$\perp$(a):
-$e_{1}$evel 1:$\overline{e_{3}}$$e_{2}$evel 2:$\overline{e_{4}}$$e_{5}$$e_{6}$evel 3:$e_{8}$$e_{11}$$\overline{x}$$\overline{e_{7}}$$e_{10}$$e_{9}$$\overline{e_{12}}$$\perp$(b):
+[…]
+[…]
 Figure 6.
@@
 This breaks a usual invariant for CDCL solvers.
-This would only be possible in a solver that permits chronological backtracking [NR18, MB19]; xMapleLCM, however, does not support this.
+This would only be possible in a solver that permits chronological backtracking [NadelRychin:chronological, MohleBiere:backtracking]; xMapleLCM, however, does not support this.
 The previous reasoning needs some clarification for the case when the extension variable $z$ with definition $z\leftrightarrow l_{1}\land l_{2}$ we want to use has already been introduced.
@@
 We have implemented the DIP-based clause learning schemes described in Section 5.2 on top of the xMapleLCM ER framework [^4] All sources used for this evaluation can be found in https://github.com/chjon/xMapleSAT/tree/main..
-We started our experimental evaluation running a variant of DIP-based learning on benchmarks of the SAT Competition [BHI+23] and comparing it to MapleLCM [LLX+17], the CDCL SAT solver on which it is based.
+We started our experimental evaluation running a variant of DIP-based learning on benchmarks of the SAT Competition [SATcomp2023] and comparing it to MapleLCM [MapleLCM], the CDCL SAT solver on which it is based.
 Even though there did not seem to be a systematic improvement on all benchmarks, a few families with important speedups were identified.
@@
 Finally, we evaluated the performance of the 1-clause variant only learns the post-DIP clause.
-After that, we report on the performance of a variety of state-of-the-art solvers, each with some distinguished characteristic: Kissat 4.0.2 [BFFH20] (an extremely efficient CDCL solver), CryptoMiniSat 5.12.1 [SNC09] (support for XOR reasoning), SBVA-CaDiCaL [HGH23] (introduction of new variables via SBVA and winner of the main track of the 2023 SAT Competition), GlucosER [AKS10] (extended-resolution based CDCL solver) and the best available configuration for xMapleLCM.
+After that, we report on the performance of a variety of state-of-the-art solvers, each with some distinguished characteristic: Kissat 4.0.2 [kissat] (an extremely efficient CDCL solver), CryptoMiniSat 5.12.1 [cryptominisat] (support for XOR reasoning), SBVA-CaDiCaL [SBVA] (introduction of new variables via SBVA and winner of the main track of the 2023 SAT Competition), GlucosER [Audemard2010ARO] (extended-resolution based CDCL solver) and the best available configuration for xMapleLCM.
 6.1. Tseitin Formulas
 These formulas have already been described in Section 5.1.
-In this section, we consider three types of unsatisfiable formulas generated by CNFgen [LENV17]: (i) Tseitin formulas on the grid, where the graph is a rectangular grid and each vertex is connected to its 4 neighbors, except for the vertices on the boundaries of the grid which have fewer neighbors, (ii) 4-regular Tseitin formulas, where the graph is a $4$-regular random one, and (iii) 6-regular Tseitin formulas built on top of $6$-regular random graphs.
+In this section, we consider three types of unsatisfiable formulas generated by CNFgen [CNFGen]: (i) Tseitin formulas on the grid, where the graph is a rectangular grid and each vertex is connected to its 4 neighbors, except for the vertices on the boundaries of the grid which have fewer neighbors, (ii) 4-regular Tseitin formulas, where the graph is a $4$-regular random one, and (iii) 6-regular Tseitin formulas built on top of $6$-regular random graphs.
 The time limit was set to 600 seconds per instance.
@@
 the size of the problem.
-However, we went one step further and studied how large the generated DRAT proofs were.
+However, we went one step further and studied how large the generated trimmed DRAT proofs were.
 On the right plot, we show the number of resolution steps in the proof.
@@
 6.2. Xorified Random k-XOR Formulas
-Another family where DIP-based systems perform very well are random $k$-xor formulas, where xorification [BS09] has been applied, i.e., replacing variables by xors of fresh variables.
+Another family where DIP-based systems perform very well are random $k$-xor formulas, where xorification [BenSasson:sizespace] has been applied, i.e., replacing variables by xors of fresh variables.
 We used CNFgen to obtain 2 sets of formulas: xor constraints of length 3 applying xorification with 2 and 3 variables.
@@
 By removing some edges from this graph, an unsatisfiable problem is generated.
-A more detailed description can be found in [BHI+23].
-We downloaded the 23 instances submitted to the SAT Competition, using the Global Benchmark Database [IS19], and observed that no system without extended-resolution reasoning could solve any benchmark in a time limit of 2 hours.
+A more detailed description can be found in [SATcomp2023].
+We downloaded the 23 instances submitted to the SAT Competition, using the Global Benchmark Database [GlobalBenchmarkDatabase], and observed that no system without extended-resolution reasoning could solve any benchmark in a time limit of 2 hours.
 We used a much larger timeout than with the two previous families because for this family we are not able to construct increasingly larger benchmarks and see how the solvers scale.
@@
 Finally, the third block compares our best xMapleLCM configuration with state-of-the-art solvers.
-It is remarkable that, apart from GlucosER, no system is able to process any instance within the given time limit.
+xMapleLCM and GlucosER are the only systems able to process any instance within the given time limit.
 Hence, the use of extended resolution is crucial in this set of benchmarks.
@@
 Otherwise, the search would be totally dominated by DIPs.
-In the left cactus plot of Figure 9 we can see that our baseline DIP-based xMapleLCM solver outperforms the CDCL-based MapleLCM that it is based on.
-When we analyzed the concrete benchmarks where our DIP-based solver outperforms MapleLCM and other CDCL-based solvers like Kissat, we observed that for all of them the percentage of decisions on extended variables is surprisingly large (over 10% of the decisions), whereas for the remaining benchmarks it is very low (in 70% of the benchmarks less than 1% of decisions are on extended variables).
-Moreover, this happens no matter how many initial variables there are, even though in problems with few number of initial variables the introduction of a few extended variables could quickly dominate the decision heuristic.
-This is remarkable since it shows that already existing decision heuristics like VSIDS or LRB somehow infer whether the newly introduced variables improve the behavior of the system.
+In the left CDF plot of Figure 10 we can see that our baseline DIP-based xMapleLCM solver outperforms the CDCL-based MapleLCM that it is based on.
+When we analyzed the concrete benchmarks where our DIP-based solver outperforms MapleLCM and other CDCL-based solvers like Kissat, we observed that for all of them the percentage of decisions on extended variables by xMapleLCM is quite large (over 10% of the decisions), whereas for the remaining benchmarks it is very low (in 70% of the benchmarks less than 1% of decisions are on extended variables).
+Moreover, this happens no matter how many initial variables there are, even though in problems with few number of initial variables the introduction of a few extended variables could potentially quickly dominate the decision heuristic.
+This is important since it shows that already existing decision heuristics like VSIDS or LRB somehow infer whether the newly introduced variables improve the behavior of the system.
 This is why we implemented, on top of our baseline DIP system, a procedure that computes the percentage of decisions on extended variables.
 If after a given number of conflicts it is still lower than 3%, it discontinues DIP learning and performs 1-UIP learning from that moment onwards.
-This is a very preliminary step in the direction of trying to automate the decision of whether to apply DIP reasoning or not, but the outcome, which is found on the left cactus plot of Figure 9, is very positive, showing that a solver disabling DIP reasoning (xMapleLCM-disabling in the plot) when the percentage of decisions on extended variables is low is still able to solve Tseitin, random $k$-xor and Interval benchmarks, but the overall performance is not significantly worsened on the remaining instances.
-Finally, the right cactus plot shows the impact of changing the type of DIP into the disabling solver: the heuristic and middle DIPs, the latter being used in the baseline configuration, are much better than random and closest ones.
+This is a very preliminary step in the direction of trying to automate the decision of whether to apply DIP reasoning or not, but the outcome, which is found on the left CDF plot of Figure 10, is very positive, showing that a solver disabling DIP reasoning (xMapleLCM-disabling in the plot) when the percentage of decisions on extended variables is low is still able to solve Tseitin, random $k$-xor and Interval benchmarks, but the overall performance is not significantly worsened on the remaining instances.
+Finally, the right CDF plot shows the impact of changing the type of DIP into the disabling solver: the heuristic and middle DIPs, the latter being used in the baseline configuration, are much better than random and closest ones.
 Figure 10. Performance of DIP variants on 2023 SAT Competition benchmarks.
@@
 This is an improvement w.r.t.
-the results in the GlucosER paper [AKS10], where a few initial benchmarks were identified but it was not proved that extended-resolution methods always outperform CDCL on increasingly more complex instances.
+the results in the GlucosER paper [Audemard2010ARO], where a few initial benchmarks were identified but it was not proved that extended-resolution methods always outperform CDCL on increasingly more complex instances.
   • We have proved that, even though the way extended variables are identified is completely different in xMapleLCM and GlucosER, they both behave well on the same instances.
-This is remarkable because our preliminary experiments revealed that randomly choosing arbitrary pairs of (non-DIP) literals to create definitions does not provide any speedup w.r.t.
+On the other hand, our preliminary experiments revealed that randomly choosing arbitrary pairs of (non-DIP) literals to create definitions does not provide any speedup w.r.t.
 CDCL.
-Hence, both methods are able to identify semantic properties about concrete literals that make them suitable for creating an extended variable definition.
+Hence, both GlucosER and xMapleLCM are able to identify meaningful semantic properties about concrete literals that make them suitable for creating an extended variable definition.
+The data we collected show that GlucosER typically introduces around four times more extension variables than xMapleLCM and that the extended variables generated by GlucosER and xMapleLCM are neither disjoint nor totally overlap.
+Only about 25% of the definitions over original variables introduced in xMapleLCM are also introduced in GlucosER.
   • We have learned that in most of the families where extended resolution works there is a large xor component.
@@
 We introduce a novel extended resolution clause learning (ERCL) algorithm that, when implemented on top of a CDCL solver, turns out to be beneficial for a variety of problems, in particular Tseitin, random $k$-xor and interval matching formulas.
-This is remarkable since automating powerful proof systems is well-known to be a difficult task.
-Somewhat surprisingly, the only previously existing attempt to incorporate ER into CDCL performs reasonably well on the same instances.
-We have hence indentified classes of formulas for which extended resolution can be automated in at least two different ways and still outperform CDCL.
-Considering the different nature of the two methods, this clearly deserves further study.
+We view this as a step towards more effective methods for incorporating extended resolution to form more powerful proof systems.
+We show that the only previously existing attempt to incorporate ER into CDCL performs reasonably well on the same instances.
+Hence, we have indentified classes of formulas for which extended resolution can be automated in at least two different ways and still outperform CDCL.
+Considering the different nature of the two methods, this deserves further study.
 Further, we also introduce a new heuristic that allows our ERCL solver xMapleSAT to perform similarly to the baseline CDCL solver MapleLCM, thus being able to get the best of both worlds, i.e., the benefit of ERCL, without sacrificing performance of the CDCL solver on, say, SAT Competition Main Track 2023 instances.
@@
 Also, the use of $k$-IPs with $k>2$ is part of our next steps.
+Finally, we intend to evaluate the impact of our technique on benchmarks from approximate hash-based counting, for which XOR reasoning is crucial.
 On the theoretical side, challenging questions like determining whether DIP- or $k$-IP based ER simulates unrestricted ER are going to be central to our research efforts.
+Acknowledgements. We thank the anonymous referees for helpful comments and corrections.